### v1

# Categorical Reparameterization with Denoising Diffusion models

Abstract
Gradient-based optimization with categorical variables typically relies on score-function estimators, which are unbiased but noisy, or on continuous relaxations that replace the discrete distribution with a smooth surrogate admitting a pathwise (reparameterized) gradient, at the cost of optimizing a biased, temperature-dependent objective. In this paper, we extend this family of relaxations by introducing a diffusion-based soft reparameterization for categorical distributions. For these distributions, the denoiser under a Gaussian noising process admits a closed form and can be computed efficiently, yielding a training-free diffusion sampler through which we can backpropagate. Our experiments show that the proposed reparameterization trick yields competitive or improved optimization performance on various benchmarks.
Machine Learning, ICML
1 Introduction
Many learning problems involve discrete choices– actions in reinforcement learning, categorical latent variables in variational inference, token-level decisions in sequence modeling, or combinatorial assignments in structured prediction and discrete optimization. A common primitive is the minimization of an objective of the form $\mathbb{E}_{\pi_{\!\theta}}[f(X)]$, where $\pi_{\!\theta}$ is a categorical distribution, often with a mean-field structure, over $L$ discrete variables, each taking values in a vocabulary of size $K$. The function $f$ represents a downstream loss or constraint penalty evaluated on discrete samples, typically through one-hot encodings. Computing $\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]$ exactly is generally intractable: in the absence of exploitable structure in $f$, it requires summing over $K^{L}$ configurations. The challenge, therefore, is to construct gradient estimators that are both computationally feasible and have a low mean squared error.
Existing estimators exhibit a standard bias–variance trade-off. Score-function estimators, such as REINFORCE (williams1992simple; greensmith2004variance), are unbiased but often suffer from high variance, which motivates the use of variance-reduction techniques most often using learned control variates (tucker2017rebar; grathwohl2018relax); they often yield useful gradients in practice but are biased with respect to the true discrete objective, with recent refinements such as ReinMax improving the approximation (liu2023reinmax). Continuous relaxations based on approximate reparameterizations, most notably the Gumbel-Softmax / Concrete construction (maddison2016concrete; jang2017categorical), replace $\pi_{\!\theta}$ by a smooth family on the simplex controlled by a temperature parameter. While this enables pathwise differentiation, taking the temperature small to reduce bias drives the sampler towards an argmax map and typically leads to ill-conditioned or vanishing gradients, whereas larger temperatures optimize a substantially different objective.
In this work, we revisit continuous relaxations through the lens of denoising diffusion models (sohl2015deep; song2019generative; ho2020denoising). Diffusion models generate data by transforming a Gaussian sample into a sample from the target data distribution through iterative denoising dynamics, which are explicitly constructed as the reverse of a chosen forward noising process. In practice, implementing the sampler requires only access to a denoiser; that is, a function that, given a noisy input and its noise level or time index, returns the expected clean signal.
Contributions.
In this paper, we exploit the key observation that for a categorical distribution supported on simplex vertices, the denoiser at each noise level can be computed in closed form. This enables us to construct a training-free, diffusion-based, differentiable, and approximate sampling map from Gaussian noise to the categorical distribution $\pi_{\!\theta}$. We then analyze the small-noise regime, which serves as the temperature parameter, and characterize the emergence of nearly constant transport regions and sharp decision boundaries, explaining when and why gradients become uninformative as the relaxation approaches the discrete target.We derive practical gradient estimators, including hard variants, that recover the hard Straight-Through (bengio2012estimating) and ReinMax (liu2023reinmax) as special cases when using a single diffusion step. We also propose a parameter-dependent initialization that improves performance while keeping the diffusion overhead small. Empirically, preliminary experiments show that our approach yields competitive or improved optimization performance on various benchmarks.
Notation.
We denote the $K$-simplex by $\Delta^{K-1}$. For a matrix $x\in{\mathbb{R}}^{L\times K}$, we write $x^{i}\in{\mathbb{R}}^{K}$ for its $i$-th row and $x^{ij}$ for the $(i,j)$-th element. The softmax operator on a matrix $x\in{\mathbb{R}}^{L\times K}$ is defined row-wise by $\text{softmax}(x)\in{\mathbb{R}}^{L\times K}$ with entries $\text{softmax}(x)^{ik}=\exp(x^{ik})/\sum_{j=1}^{K}\exp(x^{ij})$ for $(i,k)\in[L]\times[K]$. For a map $f:{\mathbb{R}}^{d}\to{\mathbb{R}}^{m}$, we write $\operatorname{J}_{x}\!f\in{\mathbb{R}}^{m\times d}$ for its Jacobian matrix. To write Jacobians for maps $f:{\mathbb{R}}^{L^{\prime}\times K^{\prime}}\to{\mathbb{R}}^{L\times K}$ conveniently, we implicitly identify matrices with their vectorized forms, obtained by stacking all rows into a single column vector. Gradients and Jacobians are taken with respect to these vectorized representations, and we do not distinguish notationally between a matrix and its vectorization.
2 Background
We consider optimization problems where the objective is an expectation with respect to a discrete distribution over a finite vocabulary $\mathsf{X}$, of the form
$F(\theta)=\mathbb{E}_{\pi_{\!\theta}}[f_{\theta}(X)]\coloneqq\sum_{x\in\mathsf{X}}f_{\theta}(x)\,\pi_{\!\theta}(x)\;,$ (1)
where $f:\mathsf{X}\times\Theta\to{\mathbb{R}}$, $\Theta\subseteq{\mathbb{R}}^{m}$, and $\{\pi_{\theta}:\theta\in\Theta\}$ is the parameterized family of probability mass functions (p.m.f.) over $\mathsf{X}$. Without loss of generality, we assume that $\mathsf{X}=\mathsf{V}^{L}$ for some $L\in\mathbb{N}$, where $\mathsf{V}\coloneqq\{e_{k}\}_{k=1}^{K}$ denotes the set of $K$ one-hot encodings, and $e_{k}$ is the one-hot vector with $1$ at position $k$.We also assume that the distribution $\pi_{\!\theta}$ factorizes according to this categorical structure: for any $\theta\in\Theta$ and $x=(x^{1},\ldots,x^{L})\in\mathsf{X}$,
$\pi_{\!\theta}(x)=\prod_{i=1}^{L}\pi^{i}_{\!\theta}(x^{i}),\;\pi^{i}_{\!\theta}(x^{i})\coloneqq\frac{\exp(\langle x^{i},\varphi_{\theta}^{i}\rangle)}{\sum_{j=1}^{K}\exp(\langle x^{j},\varphi_{\theta}^{i}\rangle)}\;,\!\!$ (2)
where $\theta\mapsto\varphi_{\theta}\in{\mathbb{R}}^{L\times K}$ is such that $\varphi^{i}_{\theta}$ are the logits of the $i$-th categorical component. The factorization (2) is standard and is used in reinforcement learning to model policies (wu2018variance; berner2019dota; vinyals2019alphastar), in training Boltzmann machines (hinton2012practical), in VQ-VAEs (van2017neural), and more recently for modeling transitions in discrete diffusion models.(hoogeboom2021argmax; austin2021structured; campbell2022continuous; lou2023discrete; shi2024simplified; sahoo2024simple).
Under mild regularity assumptions on $f$ and $(\varphi^{i}_{\theta})_{i=1}^{L}$, the gradient of (1) is given by
$\nabla_{\!{\theta}}F(\theta)=\mathbb{E}_{\pi_{\!\theta}}[\nabla_{\!{\theta}}f_{\theta}(X)]+\sum_{x}f_{\theta}(x)\nabla_{\!{\theta}}\pi_{\!\theta}(x)$ (3)
and is intractable as the sum ranges over $K^{L}$ states. Nevertheless, in the case where $f$ is separable across the dimensions, i.e., $f_{\theta}(x)=\sum_{i}^{L}f^{i}_{\theta}(x^{i})$, for a family of functions $\{f^{i}:\mathsf{V}\times\Theta\to{\mathbb{R}}\}_{i=1}^{L}$, as is the case for many information-theoretic divergences, the expectation separates into $L$ independent terms, reducing the computation to $\mathcal{O}(LK)$. Furthermore, while the first term can be approximated via Monte Carlo, the second term can also be estimated similarly using the REINFORCE identity (williams1992simple),
$\sum_{x}f_{\theta}(x)\nabla_{\!{\theta}}\pi_{\!\theta}(x)=\mathbb{E}_{\pi_{\!\theta}}[f_{\theta}(X)\nabla_{\!{\theta}}\log\pi_{\!\theta}(X)]\;.$
It is well known, however, that the vanilla Monte Carlo estimator of the r.h.s. suffers from high variance (sutton_barto_2018). In practice, it is used together with baselines or other control-variate techniques to reduce variance (greensmith2004variance; mnih2014neural; mnih2016variational; tucker2017rebar; titsias2022double; grathwohl2018relax). Other estimation methods have been proposed, such as the Straight-Through estimator (bengio2012estimating) or Gumbel-Softmax reparameterization (maddison2016concrete; jang2017categorical), which we now review. For ease of presentation and without loss generality, we assume in the remainder of the paper that $f$ does not depend on $\theta$, i.e., $f_{\theta}(x)=f(x)$.
Straight-Through and ReinMax estimators.
Popular estimators either replace the objective $F$ by a differentiable surrogate and use its gradient, or directly construct a surrogate for $\nabla_{\!{\theta}}F$ itself. From now on we assume that $f$ is differentiable w.r.t. $x$. One such estimator is the Straight-Through (ST) approach, which replaces the discrete objective by the surrogate obtained by swapping $f$ and the expectation in (1), and differentiates the map $\theta\mapsto f\!\bigl(\mathbb{E}_{\pi_{\!\theta}}[X]\bigr)$. Noting that the Jacobian of $\varphi_{\theta}\mapsto\mathbb{E}_{\pi_{\!\theta}}[X]$ is $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\in{\mathbb{R}}^{LK\times LK}$, the gradient of this surrogate w.r.t. the logits is $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\nabla_{\!{x}}f(\mathbb{E}_{\pi_{\!\theta}}[X]).$ A popular practical instance of ST replaces the expectation inside $\nabla_{\!{x}}f$ with a single Monte Carlo sample $X\sim\pi_{\!\theta}$, often referred to as hard ST;
$\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)\coloneqq\operatorname{J}_{\theta}\!\varphi_{\theta}\!{}^{\top}\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\nabla_{\!{x}}f(X)\;.$ (4)
This gradient estimator was first considered by hinton2012nnml_lec9 in the context of training with hard thresholds, where the backward pass treats the threshold operation as the identity. It was later formalized by bengio2012estimating for quantization-aware training of deep networks. The resulting gradient estimator is often effective in practice but is, by construction, biased with respect to the true discrete objective. When $f$ is linear, hard ST yields an unbiased gradient of $F$. As observed by liu2023reinmax, the hard ST surrogate can be interpreted as an unbiased estimator of a first-order approximation of $\nabla_{\!{\theta}}F(\theta)$ defined in Appendix B.
The ReinMax estimator (liu2023reinmax) improves upon hard ST by using an unbiased estimator of a second-order approximation of $\nabla_{\!{\theta}}F(\theta)$, obtained via the trapezoidal rule. In Appendix B, we prove that the resulting estimator has the following simple form, which closely resembles hard ST:
$\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)\\ \coloneqq\frac{1}{2}\operatorname{J}_{\theta}\!\varphi_{\theta}\!{}^{\top}\!\big\{\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+\widehat{C}_{\theta}(X)\big\}\nabla_{\!{x}}f(X)$ (5)
where $X\sim\pi_{\!\theta}$ and $\widehat{C}_{\theta}(X)\coloneqq(X-\mathbb{E}_{\pi_{\!\theta}}[X])(X-\mathbb{E}_{\pi_{\!\theta}}[X])^{\top}$ is an unbiased estimator of $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)$. While hard ST is exact for linear $f$, ReinMax is exact for quadratic $f$. A more detailed discussion of ReinMax is provided in Appendix B.
Continuous relaxations and soft reparameterizations.
A second family of gradient estimators relies on approximate reparameterization techniques for discrete random variables. The reparameterization trick was originally introduced for continuous distributions, where a sample can be expressed as a deterministic transformation of an auxiliary latent variable (kingma2013auto). Specifically, we temporarily assume that $\pi_{\!\theta}$ is a distribution that admits a reparameterization, that is,
$\pi_{\!\theta}\coloneqq\text{Law}\big(T_{\theta}(Z)\big),$ (6)
where $Z\sim p$ with $p$ a distribution that does not depend on $\theta$, and $T_{\theta}$ is a measurable transformation mapping $Z$ to a sample distributed according to $\pi_{\!\theta}$. Assume also that for $p$-almost every $z$ the map $\theta\mapsto T_{\theta}(z)$ is differentiable for any $\theta\in\Theta$ and that $z\mapsto\nabla_{\!{\theta}}f(T_{\theta}(z))$ satisfies standard domination conditions for all $\theta\in\Theta$, so that differentiation under the expectation is justified by the Lebesgue dominated convergence theorem. Then
$\nabla_{\!{\theta}}F(\theta)$ $=\nabla_{\!{\theta}}\mathbb{E}[f(T_{\theta}(Z))]$
$=\mathbb{E}[\operatorname{J}_{\theta}\!T_{\theta}(Z)^{\top}\nabla_{\!{x}}f(T_{\theta}(Z))]\;,$ (7)
which yields a low-variance Monte Carlo estimator of the objective gradient (schulman2015gradient). Following previous works with refer to such estimators as pathwise or reparameterized gradient.
In the discrete case however, such an exact reparameterization is not available. Indeed, any representation of $\pi_{\!\theta}$ as the pushforward of a simple continuous base distribution typically yields a map $\theta\mapsto T_{\theta}(z)$ that is piecewise constant with jump discontinuities. In particular, $\operatorname{J}_{\theta}\!T_{\theta}(z)=0$ for almost every $(z,\theta)$ and so $\mathbb{E}[\operatorname{J}_{\theta}\!T_{\theta}(Z)]=0$ while $\nabla_{\!{\theta}}F(\theta)\neq 0$, illustrating that (7) does not hold in this discrete setting because the differentiability at every $\theta\in\Theta$ breaks and the domination requirement for swapping limit and integral is violated. As a simple example, consider $\pi_{\!\theta}=\text{Bernoulli}(\theta)$ and $f(x)=x$. A possible choice of transform is $T_{\theta}(Z)=\mathds{1}_{\left(-\infty,\theta\right]}(Z)$ with $Z\sim\text{Uniform}([0,1])$ which yields to zero gradient almost surely and therefore (7) does not hold. To circumvent this problem, one typically uses continuous relaxations of $\pi_{\!\theta}$ for which (7) is valid and that trade bias for lower-variance gradients
The Gumbel–Softmax (or Concrete) distribution (maddison2016concrete; jang2017categorical) is a canonical example of such a relaxation: it replaces the categorical distribution $\pi_{\!\theta}$ (on the edges of the simplex) with a temperature-indexed family of continuous distributions $(\pi^{\theta}_{\tau})_{\tau>0}$ on the simplex, each of which admits a pathwise gradient estimator satisfying (7). Specifically, $\pi^{\theta}_{\tau}\;\coloneqq\;\text{Law}\big(T^{\theta}_{\tau}(G)\big)$ is used as a relaxed surrogate for $\pi_{\!\theta}$, where for all $\theta\in\Theta$,
$T^{\theta}_{\tau}(G)\;\coloneqq\;\text{softmax}\big((\varphi_{\theta}+G)/\tau\big)\;,\quad\tau>0$
and $G\in{\mathbb{R}}^{L\times K}$ is a random matrix with i.i.d. Gumbel entries $G^{ij}\sim\mathrm{Gumbel}(0,1)$. As $\tau\to 0$, $\{\pi^{\theta}_{\tau}:\,\tau>0\}$ converges in distribution to the distribution of the random matrix $T_{\theta}(G)\coloneqq\operatorname*{argmax}_{x\in\mathsf{X}}(\varphi_{\theta}+G)x^{\top}$ which is precisely $\pi_{\!\theta}$; see gumbel1954statistical. This is known as the Gumbel-max trick. It is easy to verify that for the surrogate objective $F_{\tau}(\theta)=\mathbb{E}_{\pi^{\theta}_{\tau}}[f(X)]$, which converges to $F$ as $\tau\to 0$, (7) holds under appropriate assumptions on $f$, thus allowing an approximate reparameterization trick at the expense of a certain bias controlled by the parameter $\tau$.
3 Method
In this section we present ReDGE (Reparameterized Diffusion Gradient Estimator), which builds upon diffusion models to define an approximate pathwise gradient for $\pi_{\!\theta}$. We start by recalling the basics of these models.
3.1 Diffusion models.
We present denoising diffusion models (DDMs) (sohl2015deep; song2019generative; ho2020denoising) and the DDIM framework (song2021ddim) through the interpolation viewpoint (liu2023flow; lipman2023flow; albergo2023stochastic). We provide more details in Appendix C. DDMs define a generative procedure for a data distribution $\pi_{0}$ by first specifying a continuous family of marginals $(\pi_{t})_{t\in[0,1]}$ that connects $\pi_{0}$ to the simple reference distribution $\pi_{1}\coloneqq\mathcal{N}(0,\mathbf{I})$. More precisely, $\pi_{t}=\text{Law}(X_{t})$, where
$X_{t}=\alpha_{t}X_{0}+\sigma_{t}X_{1}\;,\quad X_{0}\sim\pi_{0}\;,\quad X_{1}\sim\pi_{1}\;.$ (8)
Here $X_{0}$ and $X_{1}$ are independent and $(\alpha_{t})_{t\in[0,1]}$ and $(\sigma_{t})_{t\in[0,1]}$ are non-increasing and non-decreasing, respectively, schedules with boundary conditions $(\alpha_{0},\sigma_{0})\coloneqq(1,0)$ and $(\alpha_{1},\sigma_{1})\coloneqq(0,1)$. A popular example is the linear schedule, defined by $(\alpha_{t},\sigma_{t})=(1-t,t)$ (lipman2023flow; esser2024scaling). To generate new samples, DDMs simulate a time-reversed Markov chain. Given a decreasing sequence $({t_{k}})_{k=0}^{n-1}$ of $n$ time steps with ${t_{n-1}}=1$ and ${t_{0}}=0$, reverse transitions are iteratively applied to map a sample from $\pi_{{t_{k+1}}}$ to one from $\pi_{{t_{k}}}$, thereby progressively denoising until the clean data distribution $\pi_{0}$ is reached.
The DDIM framework (song2021ddim) introduces a general family of reverse transitions for denoising diffusion models. It relies on a schedule $(\eta_{t})_{t\in[0,1]}$, satisfying $\eta_{t}\leq\sigma_{t}$ for all $t\in[0,1]$, along with a family of conditional distribution given for $s<t$ by
$q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|x_{0},x_{1})\coloneqq\mathrm{N}(x_{s};\,\alpha_{s}x_{0}+\sqrt{\sigma_{s}^{2}-\eta_{s}^{2}}\,x_{1},\,\eta_{s}^{2}\mathbf{I})\;.$
When $\eta_{s}=0$, this Gaussian is understood, by abuse of notation, as a Dirac delta centered at the same mean. Clearly, for all $\eta_{s}\in[0,\sigma_{s}]$, a sample from $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\cdot|X_{0},X_{1})$ with $(X_{0},X_{1})\sim\pi_{0}\otimes\mathcal{N}(0,\mathbf{I})$ is a sample from $\pi_{s}$. We define the reverse transition
$\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{s}|x_{t})$ $=\mathbb{E}\Big[\,q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|X_{0},X_{1})\,\Big|\,X_{t}=x_{t}\Big]$ (9)
$=\mathbb{E}\Big[\,q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|X_{0},\frac{x_{t}-\alpha_{t}X_{0}}{\sigma_{t}})\,\Big|\,X_{t}=x_{t}\Big]$
where the joint distribution of the random variables $(X_{0},X_{t},X_{1})$ is defined in (8) and in the second line we have used that the $X_{1}|X_{0},X_{t}\sim\updelta_{(X_{t}-\alpha_{t}X_{0})/\sigma_{t}}$. For simplicity, we define $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(x_{s}|x_{0},x_{t})\coloneqq q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|x_{0},(x_{t}-\alpha_{t}x_{0})/\sigma_{t})$. By construction, the transitions (9) satisfy the marginalization property $\smash{\pi_{s}(x_{s})=\int\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{s}|x_{t})\,\pi_{t}(x_{t})\mathrm{d}x_{t}}$. Thus, $(\pi^{\eta}_{{t_{k}}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{k+1}}})_{k=0}^{n-2}$ defines a set of reverse transitions that enable stepwise sampling from the sequence $(\pi_{t_{k}})_{k=0}^{n-1}$. In practice, however, these transitions are intractable. A common approximation is to replace $X_{0}$ in the second line of (9) by its conditional expectations (ho2020denoising; song2021ddim). More precisely, let $\hat{x}_{0}(x_{t},t)\coloneqq\int x_{0}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$, where $\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ is defined as the conditional distribution of $X_{0}$ given $X_{t}$ in (8). Then the model proposed in (ho2020denoising; song2021ddim) corresponds to approximating each $\pi^{\eta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}$ by
$\hat{\pi}^{\eta}_{k\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k+1}(x_{k}|x_{k+1})\coloneqq q^{\eta}_{\smash{{t_{k}}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,{t_{k+1}}}}(x_{k}|\hat{x}_{0}(x_{k+1},{t_{k+1}}),x_{k+1}).$
When the denoiser $(t,x)\mapsto\hat{x}_{0}(x,t)$ is intractable it is replaced with a parametric model trained with a denoising loss.
3.2 Diffusion-based categorical reparameterization
Algorithm 1 Soft reparameterization with DDIM transitions
1:Input: grid $({t_{k}})_{k=0}^{n-1}$, schedules $(\alpha_{{t_{k}}},\sigma_{{t_{k}}},\eta_{{t_{k}}})_{k=0}^{n-1}$
2: Sample $x\sim\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$
3:for k=n1k=n-1 down to 11 do
4:  $\hat{x}_{0}\leftarrow\text{softmax}(\varphi_{\theta}+\alpha_{t_{k+1}}x/\sigma^{2}_{t_{k+1}})$
5:  $\hat{x}_{1}\leftarrow(x^{i}-\alpha_{t_{k+1}}\hat{x}_{0})/\sigma_{t_{k+1}}$
6:  $\mu\leftarrow\alpha_{{t_{k}}}\hat{x}_{0}+(\sigma_{{t_{k}}}^{2}-\eta_{{t_{k}}}^{2})^{1/2}\hat{x}_{1}$
7:  $x\leftarrow\mu+\eta_{{t_{k}}}z$   with $z\sim\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$
8:end for
9:return $x$
We now introduce our diffusion-based soft reparameterization of $\pi_{\!\theta}$. This reparameterization is built upon a DDM with target $\pi^{\theta}_{0}=\pi_{\!\theta}$. Since $\pi_{\!\theta}$ is a discrete measure, the resulting denoising distribution denoted by $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ is itself discrete. Indeed, following (8) and the factorization (2), $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\propto\prod_{i=1}^{L}\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x^{i}_{0}|x^{i}_{t})$ where
$\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x^{i}_{0}|x^{i}_{t})\propto\pi^{i}_{\!\theta}(x^{i}_{0})\mathrm{N}(x^{i}_{t};\alpha_{t}x^{i}_{0},\sigma^{2}_{t}\mathbf{I}_{K})\;.$
With this structure, the posterior-mean denoiser $\hat{x}^{\theta}_{0}(x_{t},t)\coloneqq\sum_{x_{0}}x_{0}\,\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})$ simplifies to a matrix of posterior probabilities because of the one-hot structure; i.e. we have for any $i\in[L]$ and $j\in[K]$ that $\hat{x}^{\theta}_{0}(x_{t},t)^{ij}=\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(e_{j}|x_{t})$ and the denoiser can be computed exactly and efficiently. Indeed, since $\|x^{i}_{t}-\alpha_{t}e_{j}\|^{2}=\|x^{i}_{t}\|^{2}-2\alpha_{t}x^{ij}_{t}+\alpha^{2}_{t}$, we get
$\hat{x}^{\theta}_{0}(x_{t},t)^{ij}$ $=\frac{\pi^{i}_{\!\theta}(e_{j})\exp(-\frac{\|x^{i}_{t}\|^{2}-2\alpha_{t}x^{ij}_{t}+\alpha^{2}_{t}}{2\sigma^{2}_{t}})}{\sum_{k=1}^{K}\pi^{i}_{\!\theta}(e_{k})\exp(-\frac{\|x^{i}_{t}\|^{2}-2\alpha_{t}x^{ik}_{t}+\alpha^{2}_{t}}{2\sigma^{2}_{t}})}$
$=\frac{\pi^{i}_{\!\theta}(e_{j})\exp(\alpha_{t}x^{ij}_{t}/\sigma^{2}_{t})}{\sum_{k=1}^{K}\pi^{i}_{\!\theta}(e_{k})\exp(\alpha_{t}x^{ik}_{t}/\sigma^{2}_{t}))}\;.$
This yields the following simple matrix form for the denoiser: $\hat{x}^{\theta}_{0}(x_{t},t)=\text{softmax}(\varphi_{\theta}+\alpha_{t}x_{t}/\sigma^{2}_{t})\;.\vskip 2.84544pt$ (10) Unlike standard diffusion models that learn an approximate denoiser via a neural network, here the denoiser $\hat{x}^{\theta}_{0}(\cdot,t)$ admits a closed-form expression thanks to the factorized categorical structure. This allows reverse transitions from $\pi_{1}$ to $\pi_{\!\theta}$ free of denoiser approximation and results in an approximate and differentiable sampling procedure. For simplicity we consider only the deterministic sampler corresponding to $\eta_{s}=0$ for all $s\in[0,1]$. Define
$T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{t})\coloneqq(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})\hat{x}^{\theta}_{0}(x_{t},t)+{\sigma_{s}}x_{t}/\sigma_{t}\;.$ (11)
Finally, define for all k<n2k<n-2 and $x_{1}\in{\mathbb{R}}^{L\times K}$ the DDIM mapping:
$\smash{T^{\theta}_{t_{k}}(x_{1})\coloneqq T^{\theta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\circ\dotsc\circ T^{\theta}_{t_{n-2}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{n-1}}(x_{1})}\;.$ (12)
Then, $T^{\theta}_{t_{k}}(X_{1})$ with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$ is an approximate sample from the Gaussian mixture with density $\pi^{\theta}_{\!\,{t_{k}}}(x_{t_{k}})\coloneqq\sum_{x_{0}}\prod_{i=1}^{L}\mathrm{N}(x^{i}_{t_{k}};\alpha_{t_{k}}x^{i}_{0},\sigma^{2}_{t_{k}}\mathbf{I}_{K})\pi_{\!\theta}(x_{0})$ and in particular, $T^{\theta}_{0}(X_{1})$ is the approximate and relaxed sample from $\pi_{\!\theta}$ that we use to compute the gradient estimator in (7).
Note that with a single diffusion step, the reparameterized sample is $T^{\theta}_{0}(X_{1})=\hat{x}^{\theta}_{0}(X_{1},1)=\mathbb{E}_{\pi_{\!\theta}}[X_{0}]$, because of the boundary condition $\alpha_{1}=0$, and we recover the Straight-Through estimator. On the other hand, using many diffusion steps and well-placed timesteps $({t_{k}})_{k=0}^{n-1}$ yields an almost exact reparameterization of $\pi_{\!\theta}$. As discussed previously, this is precisely what we want to avoid: the mapping becomes nearly piecewise constant in $\theta$ and we end up with a high-variance reparameterized gradient. This trade-off is directly analogous to the temperature parameter $\tau$ in Gumbel-Softmax relaxations, where a high temperature yields a relaxed but biased approximation whereas a low temperature recovers a high-variance estimator. In our case, the role of the relaxation parameter is played by the number of diffusion steps and the placement of the timesteps $({t_{k}})_{k=1}^{n-1}$.
More precisely, we show in Proposition 3.1 that the earliest timestep $t_{1}$} governs the behaviour of the reparameterized gradient: as ${t_{1}}\to 0$, the last DDIM step $T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}$ collapses almost all points in ${\mathbb{R}}^{K}$ onto a single one–hot vector, and as consequence, the Jacobian of $T^{\theta}_{0}$ w.r.t. $\theta$ vanishes. We make this intuition precise by considering next, and w.l.o.g, the case $L=1$ and $\varphi_{\theta}=\theta\in{\mathbb{R}}^{K}$. We consider the DDIM map $T^{\theta}_{0}$ as a function of the input noise $x_{1}$ and ${t_{1}}\in(0,1)$ while the remaining steps $t_{k}$} with $k\geq 2$ are fixed. For this reason we make the dependence on the timesteps $t_{1:n-1}$} explicit and write $(x_{1},{t_{1:n-1}})\mapsto T^{\theta}_{0}(x_{1};{t_{1:n-1}})$ for $T^{\theta}_{0}$. The proof is given in Appendix A.
Proposition 3.1 (Informal).
Under assumptions stated in the Appendix, and with the timesteps $({t_{k}})_{k=2}^{n-1}$ fixed, we have for all $\theta\in\Theta$, $\lim_{{t_{1}}\to 0}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1};{t_{1:n-1}})\big\|=0\;,\quad\mathbb{P}-\text{a.s.}$ (13) with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$. The proof relies on the fact that for any $t\in(0,1]$ and $x\in{\mathbb{R}}^{K}$
$\begin{cases}\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)&=\Sigma^{\theta}_{t}(x),\\ \operatorname{J}_{x}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)&=\alpha_{t}\Sigma^{\theta}_{t}(x)\big/\sigma^{2}_{t}\;,\end{cases}$ (14)
with $\Sigma^{\theta}_{t}(x)\!\coloneqq\!\mathbb{C}\mathrm{ov}_{\pi^{\theta}_{0|t}(\cdot|x)}(X_{0})$. $\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x)$ is a categorical distribution with probability vector $\text{softmax}(\theta+\alpha_{t}x/\sigma^{2}_{t})$ and under the assumption that $\alpha_{t}/\sigma^{2}_{t}\rightarrow\infty$ as $t\rightarrow 0$ it collapses into a Dirac delta unless $x$ has at least two coordinates equal to $\max_{i}x^{i}$. This leads us to consider the decision boundary $\mathsf{H}:=\bigcup_{j\neq k}\{x\in\mathbb{R}^{K}:x^{j}=x^{k}=\max_{i}x^{i}\}.$ Outside of this set, the norm of $\Sigma^{\theta}_{t}(x)$ goes to 0 when $t\to 0$ as fast as $\exp(-\alpha_{t}L(x,\theta)/\sigma^{2}_{t})$ where $L(x,\theta)>0$. Thus, both the Jacobians (14) go to 0. We then get the result by assuming that the limit, as ${t_{1}}\to 0$, of the DDIM trajectories $T^{\theta}_{{t_{1}}}(X_{1})$, with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$, land outside of $\mathsf{H}$ almost surely. We illustrate Proposition 3.1 in Figure 1.
Figure 1: Visualization of the DDIM transport for $\pi_{\!\theta}=\theta\!\cdot\!\updelta_{-2}+(1-\theta)\!\cdot\!\updelta_{2}$ with the linear schedule $(\alpha_{t},\sigma_{t})=(1-t,t)$. First two rows: DDIM trajectories with varying $t_{1}$} for two different values of $\theta\in[0,1]$. Third row: The DDIM map $\theta\mapsto T^{\theta}_{0}(x_{1};t_{1:n-1})$ for fixed input quantiles $z$ and three different values of $t_{1}$}. $\Phi$ stands for the standard Gaussian cdf.
Following the previous discussion, the timestep $t_{1}$} must be therefore be chosen in an intermediate regime, small enough to reduce the bias but not so small that the gradients become uninformative.
3.3 Hard gradient estimator.
A natural choice of hard gradient estimator is
$\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1})^{\top}\nabla_{\!{x}}f(X_{0})$ (15)
where $\smash{X_{0}\sim\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(X_{1}))}$; i.e. we draw a hard sample $X_{0}$ only at the last diffusion step.
Given that ReinMax (liu2023reinmax) provides significant improvements over the hard ST estimator, we also derive a ReinMax version of our diffusion-based reparameterization trick. First, we may re-interpret our algorithm as a composition of the reparameterization trick for continuous distribution composed with the Straight-Through gradient trick. Indeed, by the marginalization property, $\pi_{\!\theta}(x_{0})=\int\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})\,\pi^{\theta}_{{t_{1}}}(x_{t_{1}})\mathrm{d}x_{t_{1}}$, and we can write using the tower property that $\mathbb{E}_{\pi_{\!\theta}}[f(X_{0})]=\mathbb{E}_{\pi^{\theta}_{{t_{1}}}}[h_{\theta}(X_{t_{1}})]$ where $h_{\theta}(x_{t_{1}})\coloneqq\sum_{x_{0}}f(x_{0})\,\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x_{0}|x_{t_{1}})$. The Gaussian mixture $\pi^{\theta}_{{t_{1}}}$ can be reparameterized approximately using the map $T^{\theta}_{t_{1}}$ and thus we can approximate the gradient $\nabla_{\!{\theta}}\mathbb{E}_{\pi^{\theta}_{{t_{1}}}}[h_{\theta}(X_{t_{1}})]_{|\theta^{\prime}}$ with
$\nabla_{\!{\theta}}h_{\theta}(T^{\theta^{\prime}}_{{t_{1}}}(X_{t_{1}}))_{|\theta^{\prime}}+\operatorname{J}_{\theta}\!T^{\theta}_{t_{1}}(X_{1})^{\top}_{|\theta^{\prime}}\nabla_{\!{x}}h_{\theta^{\prime}}(T^{\theta^{\prime}}_{t_{1}}(X_{1}))\;.$
The only intractable terms are the gradient w.r.t. $\theta$ and $x$ of the conditional expectation $h_{\theta}$. These gradients, by abstracting away $T^{\theta^{\prime}}_{t_{1}}(X_{t_{1}})$ which is not differentiated through, are $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})$ and $\nabla_{\!{x}}h_{\theta}(x_{t_{1}})$ and are a specific case of differentiating an expectation w.r.t. the parameters of a categorical distribution, which in this case is $\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|x_{t_{1}})$. Here by using the Straight-Through approximation (4) we recover our hard gradient estimator (15); i.e. $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})\approx\nabla_{\!{\theta}}f(\hat{x}^{\theta}_{0}(x_{t_{1}},{t_{1}}))$ and $\nabla_{\!{x}}h_{\theta}(x_{t_{1}})\approx\nabla_{\!{x_{t_{1}}}}f(\hat{x}^{\theta}_{0}(x_{t_{1}},{t_{1}}))$. Our ReinMax-based estimator thus consists in using ReinMax (5) instead of hard ST as estimator for $\nabla_{\!{\theta}}h_{\theta}(x_{t_{1}})$. We coin this gradient estimator ReDGE-Max. When using a single diffusion step, i.e. ${t_{1}}=1$, $h_{\theta}$ is constant and equal to $\mathbb{E}_{\pi_{\!\theta}}[X]$ because of the boundary condition $\alpha_{1}=0$ and we recover ReinMax as a special case.
3.4 Parameter dependent $\pi_{1}$.
In the previous construction, the terminal distribution $p_{1}$ is fixed to a standard Gaussian $\pi_{1}=\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$. In our setting, however, we can exploit the factorization (2) to choose a parameter–dependent Gaussian distribution $\pi^{\theta}_{1}$ that best approximates $\pi_{\!\theta}$ in the maximum–likelihood sense. Formally, we take $\pi^{\theta}_{1}$ with factorized density $\pi^{\theta}_{1}(x)=\prod_{i=1}^{L}\mathrm{N}(x^{i};\mu^{i}_{\theta},\mathrm{Diag}(v^{i}_{\theta}))$ where for all $i\in[L]$, $(\mu^{i}_{\theta},v^{i}_{\theta})\in{\mathbb{R}}^{K}\times{\mathbb{R}}^{K}_{>0}$ and $\mathrm{Diag}(v^{i}_{\theta})\in{\mathbb{R}}^{K\times K}$ is a diagonal matrix with $v^{i}_{\theta}$ as diagonal entries. The parameters are then defined as any solution of the maximum–likelihood problem
$\{(\mu^{i}_{\theta},v^{i}_{\theta})\}_{i=1}^{L}\;\in\;\operatorname*{argmax}_{\{(\mu^{i},v^{i})\}_{i=1}^{L}}\mathbb{E}_{\pi_{\!\theta}}\bigl[\log\pi^{\theta}_{1}(X_{0})\bigr],$
where, due to the factorization of $\pi^{\theta}_{1}$, the loss writes equivalently as $\sum_{i=1}^{L}\mathbb{E}_{\pi^{i}_{\!\theta}}\bigl[\log\mathrm{N}(X_{0}^{i};\mu^{i}_{\theta},\mathrm{Diag}(v^{i}_{\theta}))\bigr].$ For each $i$, this is exactly the standard MLE problem for a multivariate Gaussian with diagonal covariance, whose one solution is given by matching the mean and per–coordinate variances of $\pi^{i}_{\!\theta}$; i.e. $\mu^{i}_{\theta}=\mathbb{E}_{\pi^{i}_{\!\theta}}[X^{i}_{0}]$ and $v^{i}_{\theta}=\mu^{i}_{\theta}\odot(1-\mu^{i}_{\theta})$. We restrict ourselves to a diagonal covariance in order to avoid expensive matrix inversions in the denoiser expression derived next. Data–dependent base distributions of this kind have also been considered in other applications, see for instance lee2022priorgrad; popov2021grad; ohayon2025posterior.
When using the base distribution $\pi^{\theta}_{1}$ and setting $\eta_{s}=0$ for all $s\in[0,1]$, the DDIM map (11) keeps the same form as before. The denoiser, however, is different and is now given in matrix form by $\hat{x}^{\theta}_{0}(x_{t},t)\!=\!\text{softmax}(\varphi_{\theta}+\frac{\alpha_{t}\lambda_{\theta}}{\sigma^{2}_{t}}\odot(x_{t}-\sigma_{t}\mu_{\theta}-\frac{\alpha_{t}}{2}\mathbf{1})).$ where $\lambda_{\theta}\in{\mathbb{R}}^{L\times K}$ with $\lambda^{i,j}_{\theta}=1/v^{i,j}_{\theta}$ and $\mathbf{1}\in{\mathbb{R}}^{L\times K}$ is the all-ones matrix. See Section C.2 for a derivation and Appendix C for the DDIM sampler with arbitrary schedule $(\eta_{s})_{s\in[0,1]}$. We refer to the resulting gradient estimator as ReDGE-Cov.
3.5 Related work
Reparameterization trick.
We discuss reparameterization tricks beyond the Gumbel-Softmax. potapczynski2020invertible replace Gumbel noise by Gaussian noise passed through an invertible transformation to obtain a more flexible family of continuous distributions on the simplex. wang2020relaxed go beyond the independence assumption (2) and propose a relaxation for correlated multivariate Bernoulli via a Gaussian copula. paulus2020gradient generalize the Gumbel–max trick by considering solutions to random linear programs and then obtain differentiable relaxations through by adding of strongly convex regularizer. Another way to obtain low variance gradient estimators is through the combination of REINFORCE with reparameterization trick-based control variates, or the use of Rao–Blackwellization (tucker2017rebar; grathwohl2018relax; liu2019rao; paulus2020raoblackwell).
Denoiser for mixture of Dirac delta.
When training a diffusion model using a dataset $(X_{i})_{i=1}^{N}$, the minimizer of the denoising loss is the denoiser for the empirical distribution $N^{-1}\!\sum_{i=1}^{N}\updelta_{X_{i}}$ and is available in closed form; see karras2022elucidating. Various recent works use this insight in different forms. scarvelis2023closed smooth the closed-form empirical denoiser to obtain training-free diffusion samplers that generalize beyond memorization. kamb2025an study denoising under architectural constraints,most notably equivariance and locality, and derive the optimal denoiser within this restricted function class, and show that the resulting training-free diffusion sampler generates novel samples and closely match the behavior of trained convolution-based diffusion models. ryzhakov2024explicit propose to train diffusion models by directly regressing to the empirical denoiser. In this work, we similarly leverage closed-form denoisers, here for distributions over $V^{L}$, but for a different goal: we use them to construct a soft reparameterization and differentiate through diffusion trajectories to obtain pathwise gradients. Finally, in concurrent work, andersson2025diffusion propose using diffusion models, in a sequential Monte Carlo setting, to generate $N$ i.i.d. reparameterized samples from the parameter-dependent empirical mixture $\smash{\sum_{i=1}^{N}w_{i}^{\theta}\,\delta_{X_{i}^{\theta}}}$, where $\smash{w_{i}^{\theta}\geq 0}$ and $\smash{\sum_{i=1}^{N}w_{i}^{\theta}=1}$, and $\theta$ denotes the state-space model parameters. This enables parameter estimation by differentiating end-to-end through the particle filter used to estimate the observation likelihood.
4 Experiments
In this section, we conduct an empirical evaluation of our method on several benchmark problems, including polynomial programming, Sudoku problem solving, and applications to variational inference and generative modeling. We compare our method against three representative baselines from the literature: the Straight-Through (ST) estimator (bengio2012estimating), the Gumbel-Softmax estimator (more precisely, its Straight-Through variant) (jang2017categorical), and the more recent ReinMax method (liu2023reinmax). Among these, ReinMax reports state-of-the-art performance on most of the benchmarks it considers and, to the best of our knowledge, is one of the most recent approaches addressing the same class of problems as ours. For this reason, and since liu2023reinmax show that ReinMax consistently outperforms several earlier alternatives, we do not include additional baselines in our comparison. All the hyperparameters are detailed in Table 5. For ReDGE and its variants we use the linear schedule $(\alpha_{t},\sigma_{t})=(t,1-t)$ and ${t_{k}}=k/(n-1)$ for $k\in[0:n-1]$ where $n$ is the number of diffusion steps. We use ReDGE-Max with $\pi_{1}=\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$ as initialization. Finally, we use the hard gradient estimator for each method.
4.1 Polynomial programming
We illustrate our approach on the polynomial programming toy problem also considered in tucker2017rebar; grathwohl2018relax; paulus2020raoblackwell; liu2023reinmax. The factorized distribution $\pi_{\!\theta}$ is such that for all $i\in[L]$, $\pi^{i}_{\!\theta}=\text{Bernoulli}(\frac{\exp(\theta^{i2})}{\exp(\theta^{i1})+\exp(\theta^{i2})})$ with $\theta\in{\mathbb{R}}^{L\times 2}$. Following prior works, we consider a fixed target vector $c=0.45$ and solve the optimization problem
$\min_{\theta\in{\mathbb{R}}^{L\times 2}}\frac{1}{L}\mathbb{E}_{\pi_{\!\theta}}\left[\|X\!\cdot\!(0\,\,\,1)^{\top}-c\!\cdot\!\mathbf{1}_{L}\|_{p}^{p}\right],$ (16)
for a fixed exponent $p\geq 1$ and with $L=128$. The optimal policy is the one that puts all the mass on the matrix with $e_{1}$ in each row. We report results on this benchmark in Fig. 2. We also emphasize several limitations of this example, which to our knowledge have not been explicitly discussed in the gradient-estimation literature and somewhat undermine its relevance as a stand-alone evaluation:
Figure 2: Polynomial programming benchmark for different values of the exponent pp.
(1).  The objective is separable and identical across dimensions, so the gradient can be recovered from only two loss evaluations (one per coordinate value).
(2).  ST estimator performs poorly in this experiment. However, note that discrete objective is determined entirely by the values of $f$ at the vertices of the product simplex. Consequently, any extension on ${\mathbb{R}}^{L\times 2}$ that matches $f$ on these vertices defines the same discrete problem, yet may induce a very different optimization landscape. As an illustration, consider the extension
$f:\ x\in\mathbb{R}^{L\times 2}\ \mapsto\ \frac{1}{L}\sum_{i=1}^{L}(-c)^{p}x^{i1}+(1-c)^{p}x^{i2}\;,$
which is linear and coincides with (16) on the vertices. For this relaxation, hard ST yields a low-variance unbiased gradient estimator (and soft ST yields the exact gradient) that performs well. See Appendix D.3 for results with this linear relaxation.
(3).  Finally, ReinMax is based on a second-order Taylor approximation of $f$. Hence, when $p=2$ in (16) the estimator is exact (proof in Appendix B.2); for other degrees it is no longer exact, though it often remains a close approximation in practice. This exactness is specific to quadratic objectives and does not extend to general $f$.
4.2 Variational Inference for Gaussian Mixture Model
As a second benchmark, we follow the Gaussian mixture variational inference experiment of liu2019rao, framed in the setting of blei2017variational.
Generative model.
We consider a $d$-dimensional Gaussian mixture model with $K$ components. In our experiments we take $d=2$, $K=20$, and draw $N=500$ observations. For $\pi\in\Delta^{K-1}$, the generative model consists in drawing i.i.d. cluster assignments $Z\coloneqq(Z^{1},\dotsc,Z^{N})\overset{\mathrm{i.i.d.}}{\sim}p_{z}\coloneqq\mathrm{Categorical}(\pi)$, then we draw a matrix $M\in{\mathbb{R}}^{K\times d}$ i.i.d. wth i.i.d. rows $M^{i}\sim p_{m}\coloneqq\mathcal{N}(0,\sigma_{0}^{2}\,\mathbf{I}_{d})$ for $i\in[K]$. Finally, we draw a matrix of $N$ observations $Y\in{\mathbb{R}}^{N\times d}$ with $Y^{i}\sim p_{y}(\cdot|M,Z^{i})\coloneqq\mathcal{N}(M^{Z^{i}},\sigma_{y}^{2}\,\mathbf{I}_{d})$.
Variational family.
Exact Bayesian inference over the means and cluster assignments $(M,Z)$ given a realization $Y=y$ is intractable, so we approximate the posterior with a mean-field variational family. Following the setup in liu2019rao, we consider the variational family
$q_{\phi}(\mathrm{d}z,\mathrm{d}m)=\pi_{\!\theta}(\mathrm{d}z)\prod_{i=1}^{K}\delta_{\hat{m}^{k}}(\mathrm{d}m^{k})\;.$
where $\phi\coloneqq(\theta,\hat{m}^{1},\dotsc,\hat{m}^{K})$ and $\pi_{\!\theta}$ is the factorized categorical distribution (2) over the cluster assignments with $L=N$ and we have implicitely replaced the cluster assignments with their one-hot encodings. Since the mean component is degenerate, rather than minimizing the KL which would require absolute continuity, we instead minimize the following objective which can be seen as a variational objective over the cluster assignment combined with a MAP estimation over the mean components;
$F(\theta;\hat{m}^{1:K})\coloneqq\sum_{i=1}^{N}\mathbb{E}_{\pi^{i}_{\!\theta}}\big[\log\pi^{i}_{\!\theta}(Z^{i})-\log p_{y}(y^{i}|\hat{m},Z^{i})\\ \textstyle-\log p_{z}(Z^{i})\big]-\sum_{k=1}^{K}\log p_{m}(\hat{m}^{k})\;.$
The results are reported in Table 1. Overall, ReDGE-Cov converges substantially faster and reaches the best final ELBO among all methods. In contrast, ReDGE-Max tracks the performance of ReinMax, while Gumbel-Softmax remains behind the vanilla ReDGE baseline both in terms of final value and speed.
We show in table 1 the final ELBO of each sampler, averaged over the 100 last iterations (with the standard deviation) as well as the final clustering accuracies.
Table 1: Gaussian Mixture Model variational inference results. We report the mean and standard deviation of the negative ELBO (NELBO) over the final 100 training iterations (lower is better), as well as the final clustering accuracy (higher is better).
Sampler Final NELBO (mean $\pm$ std) Clustering accuracy (mean $\pm$ std)
Gumbel-Softmax $1296.11\pm 87.66$ $0.56\pm 0.07$
Straight-Through $4380.48\pm 126.67$ $0.45\pm 0.04$
ReinMax $1175.73\pm 78.78$ $0.62\pm 0.10$
ReDGE $1716.58\pm 127.46$ $\mathbf{0.64\pm 0.10}$
ReDGE-Max $1186.59\pm 58.98$ $0.61\pm 0.09$
ReDGE-Cov $\mathbf{1040.04\pm 97.89}$ $0.60\pm 0.09$
4.3 Sudoku
We consider partially solved sudoku grids and frame their completion as an optimization problem. We parameterize a factorized categorical law $\pi_{\!\theta}$ over the sudoku grid with each cell represented by a categorical distribution; i.e. $\pi_{\!\theta}$ is a distribution over $\mathsf{V}^{81}$ with $V$ the set of one-hot encodings of length $9$. Each row/column/block denoted $g$ is represented by a set of indexes $i\in[81]$, and we define the digit count function $s_{g}:X\in\mathsf{V}^{81}\mapsto\sum_{i=1}^{81}\mathds{1}_{i\in g}X^{i}$, which outputs the all-ones vector exactly when the digits in $g$ form a valid permutation (i.e., each digit appears once). We use a quadratic penalty as a relaxed violation count and optimize its expectation under $\pi_{\!\theta}$: $F(\theta)\coloneqq\mathbb{E}_{\pi_{\!\theta}}[\sum_{g}\big\|s_{g}(X)-\mathbf{1}_{9}\big\|_{2}^{2}]\;.$ At first sight, taking $\pi_{\!\theta}$ to be fully factorized across cells may seem too restrictive, since valid Sudoku grids exhibit strong dependencies. The key point, however, is that while $\pi_{\!\theta}$ is mean-field conditional on a fixed $\theta$, the learning dynamics are not: the loss is highly non-separable, and each stochastic gradient step updates many cell logits jointly through shared row/column/block constraints. Consequently, dependencies are introduced through the optimization procedure itself. During training, updates are computed from random samples of the grid. Therefore the parameter iterate is itself random: after $T$ steps, $\theta_{T}$ is a random variable defined by the stochastic recursion induced by the optimizer. The distribution of an output grid produced after $T$ steps from initialization $\theta_{0}$ is thus the mixture $\mathbb{E}\!\left[\pi_{\!\theta_{T}}|\theta_{0}\right]$. Although each component in this mixture factorizes, the mixture does not: the shared optimization noise couples all coordinates through the non-separable constraints, allowing the resulting predictor to place most of its mass on globally consistent Sudoku configurations despite the mean-field parameterization.
Table 2 reports the expected quadratic constraint-violation objective (lower is better) together with the solved rate (zero violations). The diffusion-based estimators consistently improve over standard baselines: ReDGE-Cov attains the lowest mean loss, and ReDGE achieves the highest solved rate, outperforming the remaining baselines. Notably, Straight-Through yields substantially higher losses and variance, suggesting unstable optimization on this highly non-separable penalty. Among our variants, the covariance correction provides the most reliable decrease in the penalty (best mean), while ReDGE slightly trades off penalty minimization for a higher probability of reaching an exactly feasible grid (best solved %).
Table 2: Sudoku results. We report the mean and standard deviation of the loss across runs, as well as the percentage of solved Sudokus (zero violations).
Sampler Loss (mean $\pm$ std) Solved (%)
Gumbel-Softmax 14.46 $\pm$ 10.67 15.42
Straight-Through 31.29 $\pm$ 23.47 9.07
ReinMax 13.21 $\pm$ 9.47 18.31
ReDGE 9.94 $\pm$ 8.63 22.22
ReDGE-Max 11.27 $\pm$ 8.57 14.1
ReDGE-Cov 8.18 $\pm$ 7.13 20.76
4.4 Categorical VAE
Table 3: Categorical VAE results on MNIST for different latent and categorical dimensionalities. We report the final average loss (mean $\pm$ standard deviation) across runs.
Sampler L=32,K=64L=32,\,K=64 L=30,K=10L=30,\,K=10 L=48,K=2L=48,\,K=2
Gumbel-Softmax $85.37\pm 0.79$ $79.33\pm 0.50$ $88.23\pm 0.21$
Straight-Through $106.17\pm 0.03$ $101.51\pm 0.05$ $99.22\pm 0.10$
ReinMax $86.42\pm 1.50$ $80.98\pm 0.35$ $87.61\pm 0.25$
ReDGE $101.90\pm 0.49$ $89.48\pm 0.33$ $87.40\pm 0.21$
ReDGE-Max $86.46\pm 0.67$ $80.20\pm 0.68$ $87.76\pm 0.12$
ReDGE-Cov $\mathbf{81.99\pm 0.06}$ $\mathbf{78.80\pm 0.04}$ $\mathbf{82.01\pm 0.08}$
We train a Bernoulli VAE on binarized MNIST (kingma:welling:2014; rezende2015variational) following the setups in tucker2017rebar; grathwohl2018relax; liu2023reinmax. The encoder is a neural network that maps the input image $x\in\{0,1\}^{784}$ to logits $\varphi_{\theta}(x)\in{\mathbb{R}}^{L\times K}$ and defines the mean-field posterior $\pi_{\!\theta}(\cdot|x)$ (2). The decoder is a neural network modeling pixel logits $\eta_{\phi}(z)\in{\mathbb{R}}^{784}$ given a latent $z\in{\mathbb{R}}^{L\times K}$ to produce the decoding distribution $p_{\phi}(\cdot|z)=\prod_{j=1}^{784}\text{Bernoulli}\big(\sigma(\eta_{\phi}(z)^{j})\big)$ where $\sigma$ is the sigmoid function. Given a dataset $(X_{i})^{N}_{i=1}$, we optimize jointly in $(\theta,\phi)$ the negative ELBO,
$F(\theta;\phi)\coloneqq\frac{1}{N}\sum_{n=1}^{N}\mathbb{E}_{\pi_{\!\theta}(\cdot|X_{i})}\big[\log p_{\phi}(X_{i}|Z)\big]\\ -\textstyle\text{KL}\big(\pi_{\!\theta}(\cdot|X_{i})\,\big\|\,p_{z}\big)\;,$
where $p_{z}\coloneqq\text{Uniform}(\mathsf{X})$ is the discrete uniform distribution over $\mathsf{X}$. The results are summarized in Table 3 for three different configurations of $(L,K)$. ReDGE-Cov achieves a better final loss as well as a faster convergence in all the settings we considered (see more results in Section D.3), outperforming all the other baselines.
Figure 3: Categorical VAE training curves for the configuration with L=48,K=2L=48,\,K=2.
Runtime.
Our method introduces extra computation from the number of diffusion steps used. However, like Straight-Through and Gumbel-Softmax, each gradient estimate requires only a single evaluation of the objective $f$, unlike REBAR/RELAX and related variance-reduction methods which typically require multiple evaluations (tucker2017rebar; grathwohl2018relax; liu2019rao). The cost incurred by the number of diffusion steps is an $f$-independent additive overhead and is negligible in our settings where evaluating $f$ (e.g., a network forward pass or constraint computation) dominates runtime. Hence, our runtime is comparable to the competitors, since we use a small number of diffusion steps (typically $3$ to $5$). Runtime measurements are provided in Table 4.
Table 4: Average runtime per training epoch for the Categorical VAE experiment with L=48,K=2L=48,\,K=2. We report the mean and standard deviation over epochs (in seconds).
Sampler Time per epoch (s, mean $\pm$ std)
Gumbel-Softmax $5.16\pm 0.07$
Straight-Through $5.20\pm 0.20$
ReinMax $5.39\pm 0.15$
ReDGE $5.80\pm 0.32$
ReDGE-Max $6.64\pm 0.21$
ReDGE-Cov $6.15\pm 0.24$
5 Conclusion
We introduced ReDGE, a diffusion-based approach to categorical reparameterization that leverages the fact that, for categorical distributions supported on simplex vertices, the denoiser is available in closed form, yielding a training-free differentiable sampling map from Gaussian noise to $\pi_{\!\theta}$. We analyzed the small-noise regime (playing the role of a temperature) and explained how near-constant transport regions and sharp decision boundaries arise as the relaxation tightens, leading to uninformative gradients. The resulting family of estimators includes practical hard variants and recovers Straight-Through and ReinMax as one-step special cases, while allowing parameter-dependent initializations that improve performance with limited diffusion overhead.
A natural direction for future work is to explicitly correct for the residual bias. Our construction already contains all the ingredients needed for REBAR/RELAX-style control variates: a differentiable soft reparameterization, an almost exact parameter-free discrete reparameterization, and even an approximate differentiable conditional reparameterization via a forward–backward DDIM process.
Appendix A Proofs
A.1 Gradient instability: statement of Proposition 3.1 and its conditions
In this section we assume without loss of generality that $L=1$, $K\geq 2$, and $\varphi_{\theta}=\theta\in{\mathbb{R}}^{K}$. We also define for all $t\in(0,1]$, $c_{t}=\alpha_{t}/\sigma^{2}_{t}$. With these notations, noting that $\hat{x}^{\theta}_{0}(x,t)$ is the probability vector associated to $\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x)$:
$\hat{x}^{\theta}_{0}(x,t)\coloneqq\text{softmax}(\theta+c_{t}x)\;,\quad x\in{\mathbb{R}}^{K}\;,\quad\hat{x}^{\theta}_{0}(x,t)^{i}=\frac{\exp(\theta^{i}+c_{t}x^{i})}{\sum_{k=1}^{K}\exp(\theta^{k}+c_{t}x^{k})}=\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(e_{i}|x)\;,\quad i\in\{1,\ldots,K\}\;.$ (17)
In addition, recall the notation
$\Sigma^{\theta}_{t}(x)\coloneqq\mathbb{C}\mathrm{ov}(\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x))\;.$ (18)
Finally, define the union of decision boundaries:
$\mathsf{H}:=\{x\in\mathbb{R}^{K}:\text{ there exists $j,k\in[K]$, }x^{j}=x^{k}=\max_{i}x^{i}\}\;.$ (19)
We define the margin function $m:{\mathbb{R}}^{K}\to\mathbb{R}$ as the gap between the largest and second-largest coordinates
$m(x):=\max_{i}x^{i}-{m}_{2}(x)\;,\quad m_{2}(x)=\begin{cases}\max\{x^{j}\,:\,j\in\{1,\ldots,K\}\,,\,x^{j}\neq\max_{i}x^{i}\}&\text{ if there exists }x^{j}\neq\max_{i}x^{i}\\ \max_{i}x^{i}&\text{ otherwise}\;.\end{cases}$ (20)
Note that for $x\notin\mathsf{H}$, $\operatorname*{argmax}_{j\in[K]}x^{j}$ is reduced to a singleton and therefore,
$m(x):=\min_{j\neq k^{*}(x)}(x^{k^{*}(x)}-x^{j})\;,\quad k^{*}(x)=\operatorname*{argmax}_{j\in[K]}x^{j}\;.$ (21)
We now consider the following assumptions.
  • (A1)
    The schedule $(\alpha_{t},\sigma_{t})_{t\in[0,1]}$ is such that $\lim_{t\to 0}c_{t}=\infty$ where we recall that $c_{t}=\alpha_{t}/\sigma^{2}_{t}$.
Proposition A.1.
Fix $\theta\in{\mathbb{R}}^{K}$ and suppose that (A(A1)) holds. Consider the DDIM sampler $T^{\theta}_{0}:{\mathbb{R}}^{K}\to{\mathbb{R}}^{K}$ with the last time step ${t_{1}}\in(0,1)$ and all other time steps $(t_{k})_{k\geq 2}$ fixed. Then, for any $x_{1}\in{\mathbb{R}}^{K}$ such that $T^{\theta}_{t_{1}}(x_{1})\not\in\mathsf{H}$, there exists $M(t_{t})\geq 0$ only depending on $t_{2}$ such that
$\big\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})\big\|\leq 2K(K-1)(1+c_{{t_{1}}}M({t_{2}}))\exp\big(-m(T^{\theta}_{t_{1}}(x_{1}))c_{{t_{1}}}/2\big)\;.$ (22)
Consider now the additional assumption:
  • (A2)
    For any $\theta\in{\mathbb{R}}^{K}$, there exists a measurable map $\widetilde{X}^{\theta}_{0}:{\mathbb{R}}^{K}\to{\mathbb{R}}^{K}$ such that for $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$, $\mathbb{P}$-almost surely it holds
    $\lim_{{t_{1}}\to 0}T^{\theta}_{t_{1}}(X_{1})=\widetilde{X}^{\theta}_{0}(X_{1})\quad\text{ and }\quad\widetilde{X}^{\theta}_{0}(X_{1})\notin\mathsf{H}\;.$
Assumption (A(A2)) is a mild local regularity and non-degeneracy assumption on the DDIM sampler with $t_{1}$} near 0. In particular, it the number of DDIM step is equal to $1$, it easy to verify that $\lim_{{t_{1}}\to 0}T^{\theta}_{t_{1}}(X_{1})$ converges to the one-hot vector associated to $\operatorname*{argmax}_{i}X^{i}$ and therefore (A(A2)) holds. Furthermore, (A(A2)) only requires that, for each $\theta\in{\mathbb{R}}^{K}$, the trajectory $t_{1}\mapsto T^{\theta}_{t_{1}}(X_{1})$, started from Gaussian noise $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$, admits an almost-sure limit as $t_{1}\to 0$, and that this limit does not lie on the decision boundary $\mathsf{H}$. In particular, we do not assume that $\widetilde{X}^{\theta}_{0}(X_{1})$ coincides with the data distribution or that it is one-hot; we only use that the limiting state is well-defined and is not in $\mathsf{H}$.
Corollary A.2.
Fix $\theta\in{\mathbb{R}}^{K}$ and suppose that (A(A1))-(A(A2)) hold. Let $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$ and consider the DDIM sampler $T^{\theta}_{0}:{\mathbb{R}}^{K}\to{\mathbb{R}}^{K}$ with the last time step ${t_{1}}\in(0,1)$ and all other time steps $(t_{k})_{k\geq 2}$ fixed. Then, $\mathbb{P}$-almost surely
$\lim_{{t_{1}}\to 0}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1})\big\|=0\;.$ (23)
In the next section, we state and prove preliminary results needed for the proof of Proposition A.1 postponed to Section A.3.
A.2 Supporting Lemmas for Proposition A.2
Lemma A.3.
For each $t\in(0,1]$ and $x\in\mathbb{R}^{K}$,
$\operatorname{J}_{\theta}\!\hat{x}^{\theta}_{0}(x,t)=\Sigma^{\theta}_{t}(x)\;,\quad\operatorname{J}_{x}\!\hat{x}^{\theta}_{0}(x,t)=c_{t}\Sigma^{\theta}_{t}(x)\;.$
Proof.
By (17), a direct computation gives, for all $i,j\in[K]$, $x,\theta$,
$\partial_{\theta^{j}}\hat{x}^{\theta}_{0}(x,t)^{i}=\hat{x}^{\theta}_{0}(x,t)^{i}(\delta_{ij}-\hat{x}^{\theta}_{0}(x,t)^{j})\;,$
so in matrix form
$\operatorname{J}_{\theta}\!\hat{x}^{\theta}_{0}(x,t)=\mathrm{Diag}\bigl(\hat{x}^{\theta}_{0}(x,t)\bigr)-\hat{x}^{\theta}_{0}(x,t)\hat{x}^{\theta}_{0}(x,t)^{\top}\;.$
By definition, $\Sigma^{\theta}_{t}(x)=\mathbb{E}_{\pi^{\theta}_{0|t}(\cdot|x)}[X_{0}X^{\top}_{0}]-\hat{x}^{\theta}_{0}(x,t)\hat{x}^{\theta}_{0}(x,t)^{\top}$, where $(X_{0},X_{t})$ follows the distribution with density $\pi_{\!\theta}(x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0},\sigma^{2}_{t}\mathbf{I}_{K})$, and by (17)
$\mathbb{E}_{\pi^{\theta}_{0|t}(\cdot|x)}[X_{0}X^{\top}_{0}]=\sum_{i=1}^{K}e_{i}e_{i}^{\top}\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(e_{i}|x)=\mathrm{Diag}(\hat{x}^{\theta}_{0}(x,t))$
and hence the equality $\operatorname{J}_{\theta}\!\hat{x}^{\theta}_{0}(x,t)=\Sigma^{\theta}_{t}(x)$. The Jacobian w.r.t. $x$ follows using similar arguments. ∎
Lemma A.4 (Continuity of the margin function outside of $\mathsf{H}$ (19)).
$m_{2}$ is continuous on ${\mathbb{R}}^{K}\setminus\mathsf{H}$ and therefore $m$ as well.
Proof.
Note that ${\mathbb{R}}^{K}\setminus\mathsf{H}$ is the disjoint union of the open sets $\mathsf{U}_{i}=\{x\in\mathsf{H}\,:\,i=\operatorname*{argmax}_{j}x^{j}\}$. Since on $\mathsf{U}_{i}$, $m_{2}(x)=\max_{j\neq i}x^{j}$, we obtain that $m_{2}$ is continuous on ${\mathbb{R}}^{K}\setminus\mathsf{H}$. ∎
Lemma A.5 (Softmax bound).
Let $z\notin\mathsf{H}$ where $\mathsf{H}$ is defined in (19) and $p(z)\coloneqq\mathrm{softmax}(z)$. Then,
$1-p(z)^{k^{*}(z)}\leq(K-1)\,\exp\big(-m(z))\;,$ (24)
and for all $j\neq k^{*}(z)$,
$p(z)^{j}\;\leq\;\exp\big(-m(z)\big)\;.$ (25)
Proof.
For ease of notation, we simply denote $p(z)$ by $p$. Since $z\notin\mathsf{H}$, We have that
$p^{j}=\frac{\exp(z^{j})}{\sum_{\ell=1}^{K}\exp(z^{\ell})}=\frac{\exp(z^{j}-z^{k^{*}(z)})}{1+\sum_{\ell\neq k^{*}(z)}\exp(z^{\ell}-z^{k^{*}(z)})}$
and for every $j\neq k^{\star}(z)$, we have $z^{k^{\star}(z)}-z^{j}\geq m(z)$, so $z^{j}-z^{k^{\star}(z)}\leq-m(z)$ and $p^{j}\leq\exp(-m(z))$. Then
$1-p^{k^{*}(z)}=\sum_{j\neq k^{*}(z)}p^{j}\leq(K-1)\exp(-m(z))\;.$
Lemma A.6 (Covariance control).
Let $p\in\Delta^{K-1}$ and $\Sigma=\mathrm{Diag}(p)-pp^{\top}$. Let $p^{\max}:=\max_{j\in[K]}p^{j}$. Then it holds that
$\sum_{j,k=1}^{K}|\Sigma^{jk}|\;\leq\;2K\,(1-p_{\max})\;.$
As a consequence, $\|\Sigma\|\leq 2K(1-p_{\max})$, where $\|\cdot\|$ is the operator norm.
Proof.
By definition of the covariance matrix $\Sigma$, we have that $\Sigma^{jj}=p^{j}(1-p^{j})$ and $|\Sigma^{jk}|=p^{j}p^{k}$. Let $k^{*}=\operatorname*{argmax}_{i\in[K]}p^{i}$ and define $\smash{p^{\text{max}}=p^{k^{*}}}$. For all $j\in[K]$,
$\sum_{k=1}^{K}|\Sigma^{jk}|=\Sigma^{jj}+\sum_{k\neq j}p^{j}p^{k}=p^{j}(1-p^{j})+p^{j}\sum_{k\neq j}p^{k}=2p^{j}(1-p^{j})\;.$
Next, we have that $p^{j}(1-p^{j})\leq 1-p^{\text{max}}$ since if $j=k^{*}$ then $p^{j}(1-p^{j})\leq 1-p^{\text{max}}$ and if $j\neq k^{*}$ then $p^{j}(1-p^{j})\leq p^{j}\leq\sum_{\ell\neq k^{*}}p^{\ell}=1-p^{\text{max}}$. Hence
$\sum_{k=1}^{K}|\Sigma^{jk}|\leq 2(1-p^{\text{max}})\;.$
The final bound is an easy consequence of the norm equivalent in finite dimension. ∎
We define the notation $a(s,t)=\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t}$ and $b(s,t)=\sigma_{s}/\sigma_{t}$ so that the one-step map writes
$T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)=a(s,t)\,\hat{x}^{\theta}_{0}(x,t)+b(s,t)x\;.$ (26)
Lemma A.7 (DDIM Jacobian bound).
There exists a finite constant $M(t_{2})<\infty$, depending only on $t_{2}$, $K$ and the schedule $(\alpha_{t},\sigma_{t})$, such that for all $x_{1}\in{\mathbb{R}}^{K}$ and all $t_{1}\in(0,t_{2})$,
$\big\|\operatorname{J}_{\theta}\!T^{\theta}_{t_{1}}(x_{1})\big\|\;\leq\;M(t_{2})\;.$ (27)
In particular, the bound in (27) does not depend on $t_{1}$.
Proof.
Single-step bound. We start with a single-step bound on the Jacobian of the map $T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ with $s<t$. For fixed $t\geq t_{2}$ and $s\in[0,t]$, the reverse step $T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ has the form (26), so using Lemma A.3 so we obtain
$\operatorname{J}_{\theta}\!T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)$ $=a(s,t)\,\Sigma^{\theta}_{t}(x)\;,$
$\operatorname{J}_{x}\!T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)$ $=a(s,t)\,c_{t}\,\Sigma^{\theta}_{t}(x)+b(s,t)I_{K}\;.$
Since the schedule $t\mapsto(\alpha_{t},\sigma_{t},1/\sigma_{t})$ is continuous on $\left[t_{2},1\right]$, since $t_{2}>0$, the coefficients $a(s,t),b(s,t)$ and $c_{t}$ are bounded on the compact set $\{(s,t):0\leq s\leq t,\,t_{2}\leq t\leq 1\}$. Therefore, the uniform covariance bound from Lemma A.6 implies that there exist finite constants $L_{1}(t_{2}),L_{2}(t_{2})$ such that for all $t\in[t_{2},1]$, $s\in[0,t]$, $x\in{\mathbb{R}}^{K}$ and $\theta\in{\mathbb{R}}^{K}$,
$\big\|\operatorname{J}_{\theta}\!T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)\big\|\leq L_{1}(t_{2}),\qquad\big\|\operatorname{J}_{x}\!T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)\big\|\leq L_{2}(t_{2})\;.$ (28)
Bound via induction. Next, for each $k\in[1:n-1]$ we use the following notation for the parameter Jacobian
$G_{k}(x_{1},\theta_{0}):=\operatorname{J}_{\theta}\!T^{\theta}_{t_{k}}(x_{1})_{|\theta_{0}}\;.$
By construction, the initial state at time $t_{n-1}=1$ does not depend on $\theta$, so $G_{n-1}(x_{1},\theta_{0})=0$ for all $x_{1}$.
For $k=2,\dots,n-1$ we have, by definition of the sampler,
$T^{\theta}_{t_{k}}(x_{1})=T^{\theta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta}_{t_{k+1}}(x_{1})\big)\;.$
Applying the chain rule with respect to $\theta$ at $\theta_{0}$ gives
$G_{k}(x_{1},\theta)=\operatorname{J}_{\theta}\!T^{\theta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta_{0}}_{t_{k+1}}(x_{1})\big)_{|\theta_{0}}+\operatorname{J}_{x}\!T^{\theta_{0}}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\big(T^{\theta_{0}}_{t_{k+1}}(x_{1})\big)\cdot G_{k+1}(x_{1},\theta_{0})\;.$
We now show by induction that for all $k\in[2:n-1]$, there exists a constant $M_{k}({t_{2}})$ depending only on $L_{1}({t_{2}}),L_{2}({t_{2}})$ and the number of DDIM steps such that $\|G_{k}(x_{1},\theta_{0})\|\leq M_{k}({t_{2}})$. First, the constant bounding $\|G_{n-1}(x_{1},\theta_{0})\|$ is trivial. Assume then that $\|G_{k+1}(x_{1},\theta_{0})\|\leq M_{k+1}({t_{2}})$. Taking norms and applying the inequality (28) with $t=t_{k+1}\geq t_{2}$ and $s=t_{k}$ yields
$\|G_{k}(x_{1},\theta_{0})\|\leq L_{1}({t_{2}})+L_{2}({t_{2}})\,\|G_{k+1}(x_{1},\theta_{0})\|\;.$ (29)
and thus $\|G_{k}(x_{1},\theta_{0})\|\leq M_{k}({t_{2}})\coloneqq L_{1}({t_{2}})+L_{2}({t_{2}})M_{k+1}({t_{2}})$, which shows the result. ∎
A.3 Proof of the main results
Proof of Proposition A.1.
Step 1: Jacobian bounds on a compact set. Let $x\not\in\mathsf{H}$, and recall the margin function writes $m(x)=\min_{j\neq k^{*}(x)}(x^{k^{*}(x)}-x^{j})$ with $k^{*}(x)\coloneqq\operatorname*{argmax}_{j\in[K]}x^{j}$. By definition of $\mathsf{H}$, it holds then that $m(x)>0$.
Now consider the logit margin defined for all $j\neq k^{*}(x)$ by $\Delta^{j}_{t}(x,\theta)\coloneqq(\theta^{k^{*}(x)}-\theta^{j})+c_{t}(x^{k^{*}(x)}-x^{j})$. Then, letting $B(\theta)\coloneqq\max_{(i,j)\in[K]^{2}}|\theta^{i}-\theta^{j}|$, we have that
$\Delta^{j}_{t}(x,\theta)\geq-B(\theta)+c_{t}m(x)\;.$
Since $\lim_{t\to 0}c_{t}=\infty$ by (A(A1)), there exists $t_{\star}(\theta,x)$ such that for all $t<t_{\star}(\theta,x)$, $\Delta^{j}_{t}(x,\theta)\geq c_{t}m(x)/2$ and thus
$\min_{j\neq k^{*}(x)}\Delta^{j}_{t}(x,\theta)=m(\theta+c_{t}x)\geq c_{t}m(x)/2\;,$
where we have used that $k^{*}(\theta+c_{t}x)=k^{*}(x)$ since $\Delta^{j}_{t}(x,\theta)>0$ for all $j\neq k^{*}(x)$. Now define for ${t_{1}}<t_{\star}(\theta,x)$, $p^{\text{max}}(x,{t_{1}})=\max_{j\in[K]}\hat{x}^{\theta}_{0}(x,{t_{1}})^{j}$ and we recall that $\hat{x}^{\theta}_{0}(x,{t_{1}})\coloneqq\text{softmax}(\theta+c_{t_{1}}x)\in\Delta^{K-1}$. Applying Lemma A.5 with $z=\theta+c_{t_{1}}x$, we obtain
$1-p^{\text{max}}(x,{t_{1}})\leq(K-1)\exp\big(\!-\!m(\theta+c_{t_{1}}x)\big)\leq(K-1)\exp\Big(-\frac{m(x)}{2}\,c_{t_{1}}\Big)\;.$
Hence by Lemma A.6, for the covariance (18) we have that
$\big\|\Sigma_{t_{1}}(x)\big\|\leq 2K\big(1-p^{\text{max}}(x,{t_{1}})\big)\leq 2K(K-1)\exp\Big(-\frac{m(x)}{2}\,c_{t_{1}}\Big)\;.$
Using the gradient identities in Lemma A.3 $\operatorname{J}_{x}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)=c_{t_{1}}\Sigma^{\theta}_{t_{1}}(x)$ and $\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)=\Sigma^{\theta}_{t_{1}}(x)$ then for ${t_{1}}\in(0,t_{\star}(\theta,x))$, we have the following bounds
$\big\|\operatorname{J}_{x}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)\big\|$ $\leq c_{t_{1}}M_{K}\exp(-m(x)c_{t_{1}}/2)\;,$ (30)
$\big\|\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(x)\big\|$ $\leq M_{K}\exp(-m(x)c_{t_{1}}/2),$ (31)
with $M_{K}:=2K(K-1)$.
Step 2: chain rule for the parameter gradient.
For any $x_{1}\in{\mathbb{R}}^{K}$, $T^{\theta}_{0}(x_{1})=T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta}_{{t_{1}}}(x_{1})\big)$ and thus for any $\theta_{0}\in{\mathbb{R}}^{K}$,
$\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})_{|\theta_{0}}=\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)_{|\theta_{0}}\\ +\operatorname{J}_{x}\!T^{\theta_{0}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)\cdot\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1})_{|\theta_{0}}\;.$
Hence, taking the norms, we get
$\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})_{|\theta_{0}}\|$ $\leq\|\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)_{|\theta_{0}}\|+\|\operatorname{J}_{x}\!T^{\theta_{0}}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}\big(T^{\theta_{0}}_{{t_{1}}}(x_{1})\big)\|\|\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1})_{|\theta_{0}}\|$
By Lemma A.7, there exists a finite constant $M(t_{2})$ (depending only on $t_{2},K$ and the schedule) such that
$\sup_{{t_{1}}\in(0,t_{2})}\sup_{x_{1}\in{\mathbb{R}}^{K}}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{{t_{1}}}(x_{1})\big|_{\theta_{0}}\big\|\leq M(t_{2})\;.$
Finally, since by assumptions $x_{1}\in{\mathbb{R}}^{K}$ is such that $T^{\theta}_{t_{1}}(x_{1})\notin\mathsf{H}$, we get by applying the bounds (30) and (31)
$\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(x_{1})_{|\theta_{0}}\|\leq(1+c_{{t_{1}}}M({t_{2}}))M_{K}\exp\big(-m(T^{\theta_{0}}_{t_{1}}(x_{1}))c_{{t_{1}}}/2\big)\;.$
which yields the result. ∎
Proof of Proposition A.2.
The proof is an immediate consequence of Lemma A.4 and Proposition A.1. ∎
Appendix B On ReinMax
B.1 An alternative view of ReinMax
For the sake of completeness we derive the ReinMax gradient estimator from first principles and arrive at the alternative and simpler expression (5). We assume for the sake of simplicity that $L=1$ and $\varphi_{\theta}=\theta\in{\mathbb{R}}^{K}$. We restate some of the arguments in the original paper with our notation and interpretation.
First, the ground-truth gradient we seek to estimate is
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]\coloneqq\sum_{i=1}^{K}f(e_{i})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{i})\;.$ (32)
Upon baseline substraction (here $\mathbb{E}_{\pi_{\!\theta}}[X]$), we have
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\sum_{i,j}(f(e_{i})-f(e_{j}))\nabla\pi_{\!\theta}(e_{i})\,\pi_{\!\theta}(e_{j})\;.$
From this, the authors derive the first-order approximation interpretation of the ST estimator. Indeed, using a first order approximation, we get that
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]\approx\sum_{i,j}\nabla_{\!{x}}f(e_{j})^{\top}(e_{i}-e_{j})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{i})\,\pi_{\!\theta}(e_{j})\;.$
and it can be shown that the expectation of the ST gradient estimator is the r.h.s. term. Indeed, recall that the ST estimator is given by
$\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)=\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X]^{\top}\nabla_{\!{x}}f(X)\;.$ (33)
Lemma B.1.
It holds that
$\mathbb{E}_{\pi_{\!\theta}}\left[\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)\right]=\sum_{i,j}\nabla_{\!{x}}f(e_{i})^{\top}(e_{j}-e_{i})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})\,\pi_{\!\theta}(e_{i})\;.$
Proof.
We have
$\mathbb{E}\big[\widehat{\nabla}^{\text{ST}}_{\!\theta}F(X;\theta)\big]$ $=\sum_{i=1}^{K}\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[X]^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\left\{e_{j}\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})^{\top}\right\}^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})e_{j}^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})(e_{j}-e_{i})^{\top}\nabla_{\!{x}}f(e_{i})\pi_{\!\theta}(e_{i})$
$=\sum_{i,j}\nabla_{\!{x}}f(e_{i})^{\top}(e_{j}-e_{i})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{j})\pi_{\!\theta}(e_{i})$
where the penultimate line is obtained by baseline substraction. ∎
Now consider the second-order approximation of (32) obtained via Heun’s method,
$\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)]\coloneqq\sum_{i,j}\frac{1}{2}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{i})\big)^{\top}(e_{i}-e_{j})\nabla_{\!{\theta}}\pi_{\!\theta}(e_{i})\,\pi_{\!\theta}(e_{j})\;.$ (34)
We shall now obtain the ReinMax estimator by deriving an alternative expression of (34). First, note that for $(i,k)\in[K]^{2}$, $\partial_{\theta^{k}}\pi_{\!\theta}(e_{i})=\pi_{\!\theta}(e_{i})(\delta_{ik}-\pi_{\!\theta}(e_{k}))$ where $\delta_{ik}$ is the Kroenecker symbol. We thus get
$(\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)])^{k}$ $=\frac{1}{2}\sum_{i,j}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{i})\big)^{\top}(e_{i}-e_{j})\pi_{\!\theta}(e_{i})(\delta_{ik}-\pi_{\!\theta}(e_{k}))\,\pi_{\!\theta}(e_{j})$
$=\frac{1}{2}\sum_{j=1}^{K}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{k})\big)^{\top}(e_{k}-e_{j})\,\pi_{\!\theta}(e_{k})\pi_{\!\theta}(e_{j})\;.$ (35)
where we have used that $\sum_{i,j}\big(\nabla_{\!{x}}f(e_{j})+\nabla_{\!{x}}f(e_{i})\big)^{\top}(e_{i}-e_{j})\,\pi_{\!\theta}(e_{i})\pi_{\!\theta}(e_{j})=0$.
To avoid evaluating $\nabla_{\!{x}}f$ more than once, ReinMax leverages the following identity to derive the estimator:
$(\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)])^{k}$ $=\frac{1}{2}\mathbb{E}_{\pi_{\!\theta}}[\pi_{\!\theta}(e_{k})\nabla_{\!{x}}f(X)^{\top}(e_{k}-X)]+\frac{1}{2}\mathbb{E}_{\pi_{\!\theta}}\big[\langle X,e_{k}\rangle\nabla_{\!{x}}f(X)^{\top}(X-\mathbb{E}_{\pi_{\!\theta}}[X])\big]$
$=\frac{1}{2}\mathbb{E}_{\pi_{\!\theta}}\left[\nabla_{\!{x}}f(X)^{\top}\left\{\pi_{\!\theta}(e_{k})(e_{k}-X)+\langle X,e_{k}\rangle(X-\mathbb{E}_{\pi_{\!\theta}}[X])\right\}\right]$
$=\mathbb{E}_{\pi_{\!\theta}}\left[\nabla_{\!{x}}f(X)^{\top}\left\{2\frac{\pi_{\!\theta}(e_{k})+\langle X,e_{k}\rangle}{2}\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\frac{\pi_{\!\theta}(e_{i})+\langle X,e_{i}\rangle}{2}\big)-\frac{\pi_{\!\theta}(k)}{2}\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\pi_{\!\theta}(e_{i})\big)\right\}\right]\;.$
Recalling that $\partial_{\theta^{k}}\pi_{\!\theta}(e_{i})=\pi_{\!\theta}(e_{i})(\delta_{ik}-\pi_{\!\theta}(e_{k}))$, we find that
$\pi_{\!\theta}(k)\big(e_{k}-\sum_{i=1}^{K}e_{i}\,\pi_{\!\theta}(e_{i})\big)=(\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X])^{k}$
and thus that $\operatorname{J}_{\theta}\!\mathbb{E}_{\pi_{\!\theta}}[X]=\mathrm{Diag}(\mathbb{E}_{\pi_{\!\theta}}[X])-\mathbb{E}_{\pi_{\!\theta}}[X]\mathbb{E}_{\pi_{\!\theta}}[X]^{\top}$. Finally, defining the conditional distribution $\pi_{\!\theta}(\cdot|x)$ with $\pi_{\!\theta}(e_{i}|x)\coloneqq(\pi_{\!\theta}(e_{i})+\langle x,e_{i}\rangle)/2$, we see that the ReinMax estimator, defined as
$\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)\coloneqq\big[2\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}(\cdot|X)}(\widetilde{X})-\frac{1}{2}\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\big]\nabla_{\!{x}}f(X)\;,$ (36)
and that it satifies $\hat{\nabla}^{\text{2nd}}_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]$.
Remark B.2 (On the conditional distribution).
We have that $\pi_{\!\theta}(e_{i})=\sum_{j=1}^{K}\pi_{\!\theta}(e_{i}|e_{j})\pi_{\!\theta}(e_{j}).$
Finally, note that $\pi_{\!\theta}(\cdot|x)=\frac{1}{2}(\pi_{\!\theta}+\updelta_{x})$; i.e. a mixture of $\pi_{\!\theta}$ and the point mass $\updelta_{x}$. Therefore using Lemma B.3, we have
$\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}(\cdot|x)}(\widetilde{X})=\frac{1}{2}\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+\frac{1}{4}(x-\mathbb{E}_{\pi_{\!\theta}}[X])(x-\mathbb{E}_{\pi_{\!\theta}}[X])^{\top}\;,$
and plugging in (36) we recover (5); i.e.
$\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)=\frac{1}{2}\big\{\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)+(X-\mathbb{E}_{\pi_{\!\theta}}[X])(X-\mathbb{E}_{\pi_{\!\theta}}[X])^{\top}\big\}\nabla_{\!{x}}f(X)$ (37)
Lemma B.3.
Consider the following mixture $P=\frac{1}{2}P_{1}+\frac{1}{2}P_{2}$ on $\mathbb{R}^{d}$. Denote by $\mu_{i}=\mathbb{E}_{P_{i}}[X]$ and $\Sigma_{i}=\mathrm{Cov}_{P_{i}}(X)$ the means and covariances of the components $i=1,2$. The mean $\mu$ and covariance $\Sigma$ of the mixture are
$\mu=\frac{\mu_{1}+\mu_{2}}{2},\qquad\Sigma=\frac{\Sigma_{1}+\Sigma_{2}}{2}+\frac{1}{4}(\mu_{1}-\mu_{2})(\mu_{1}-\mu_{2})^{\top}.$
Proof.
Let $Z$ be the component indicator with $\mathbb{P}(Z=1)=\mathbb{P}(Z=2)=\tfrac{1}{2}$. Then
$\mathbb{E}[X\mid Z=i]=\mu_{i},\qquad\mathbb{C}\mathrm{ov}(X\mid Z=i)=\Sigma_{i},\quad i=1,2.$
By the law of total covariance,
$\Sigma=\mathbb{C}\mathrm{ov}(X)=\mathbb{E}\big[\mathbb{C}\mathrm{ov}(X\mid Z)\big]+\mathbb{C}\mathrm{ov}\big(\mathbb{E}[X\mid Z]\big).$
The first term equals $\tfrac{1}{2}(\Sigma_{1}+\Sigma_{2})$. For the second term,
$\mathbb{C}\mathrm{ov}\big(\mathbb{E}[X\mid Z]\big)=\mathbb{C}\mathrm{ov}(\mu_{Z})=\mathbb{E}[\mu_{Z}\mu_{Z}^{\top}]-\mu\mu^{\top}=\tfrac{1}{2}(\mu_{1}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top})-\mu\mu^{\top}.$
With $\mu=\frac{\mu_{1}+\mu_{2}}{2}$ a short algebraic simplification gives
$\frac{1}{2}(\mu_{1}\mu_{1}^{\top}+\mu_{2}\mu_{2}^{\top})-\mu\mu^{\top}=\frac{1}{4}(\mu_{1}-\mu_{2})(\mu_{1}-\mu_{2})^{\top},$
and combining the two terms yields the stated formula. ∎
B.2 Exactness of ReinMax for quadratic functions
In this section we show that (37) is unbiased for quadratic functions. Let $f(x)=x^{\top}Ax+b^{\top}x+d,$ with $A$ a symmetric matrix. Then $\nabla_{\!{x}}f(x)=2Ax+b$ is affine. Moreover,
$\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{C}\mathrm{ov}(\pi_{\!\theta})\,\overline{f},$
where $\overline{f}=(f(e_{i}),\dots,f(e_{K}))^{\top}$ and we used the identity $\nabla_{\!{\theta}}\mathbb{E}_{\pi_{\!\theta}}=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)$. Since $f(e_{i})=A_{ii}+b_{i}+d$, denoting $\overline{f}=\mathrm{diag}(A)+b+d\mathbf{1}$ where $\mathrm{diag}(A)\in{\mathbb{R}}^{K}$ is the vector of diagonal entries of $A$,
$\nabla_{\theta}\mathbb{E}_{\pi_{\!\theta}}[f(X)]=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\big(\mathrm{diag}(A)+b\big),$ (38)
where we have used that $\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)\mathbf{1}=0$. We now compute exactly $\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]$. Write $C_{\theta}:=\mathbb{C}\mathrm{ov}_{\pi_{\!\theta}}(X)$ and $\mu_{\theta}\coloneqq\mathbb{E}_{\pi_{\!\theta}}[X]$. Using the definition (37),
$\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]$ $=\frac{1}{2}\Big(C_{\theta}\,\mathbb{E}_{\pi_{\!\theta}}[2AX+b]+\mathbb{E}_{\pi_{\!\theta}}\big[(X-\mu_{\theta})(X-\mu_{\theta})^{\top}(2AX+b)\big]\Big)$
$=\frac{1}{2}\Big(C_{\theta}(2A\mu_{\theta}+b)+2\,\mathbb{E}_{\pi_{\!\theta}}\big[(X-\mu_{\theta})(X-\mu_{\theta})^{\top}AX\big]+C_{\theta}b\Big).$ (39)
It remains to evaluate $T:=\mathbb{E}_{\pi_{\!\theta}}[(X-\mu_{\theta})(X-\mu_{\theta})^{\top}AX]$. We have that
$T=\sum_{k=1}^{K}\pi_{\!\theta}(e_{k})\,(e_{k}-\mu_{\theta})(e_{k}-\mu_{\theta})^{\top}Ae_{k}.$
Using $e_{k}^{\top}Ae_{k}=A_{kk}$ and $\mu^{\top}Ae_{k}=(A\mu)_{k}$, expand
$(e_{k}-\mu)(e_{k}-\mu)^{\top}Ae_{k}=A_{kk}e_{k}-(A\mu)_{k}e_{k}-A_{kk}\mu+(A\mu)_{k}\mu.$
Summing over $k$ yields
$T=\big(\mathrm{diag}(\mu)-\mu\mu^{\top}\big)\mathrm{diag}(A)-\big(\mathrm{diag}(\mu)-\mu\mu^{\top}\big)(A\mu)=C\,\mathrm{diag}(A)-C(A\mu).$ (40)
Plugging (40) into (39) gives
$\mathbb{E}_{\pi_{\!\theta}}[\widehat{\nabla}^{\text{RM}}_{\!\theta}F(X;\theta)]=\frac{1}{2}\Big(C(2A\mu+b)+2(C\mathrm{diag}(A)-C(A\mu))+Cb\Big)=C\big(\mathrm{diag}(A)+b\big)\;,$
which proves the result
Appendix C DDIM with a general Gaussian reference $\pi_{1}$
C.1 Reverse transitions
Let $\pi_{0}$ be a probability distribution on ${\mathbb{R}}^{d}$. Consider the distribution path $(\pi_{t})_{t\in[0,1]}$ defined by $\pi_{t}=\text{Law}(X_{t})$, where
$X_{t}=\alpha_{t}X_{0}+\sigma_{t}X_{1},\qquad(X_{0},X_{1})\sim\pi_{0}\otimes\pi_{1}\;,$ (41)
where we assume the more general reference $\pi_{1}=\mathcal{N}(\mu,\Sigma)$ with $\Sigma\in\mathcal{S}_{++}({\mathbb{R}}^{d})$. Let $(\eta_{t})_{t\in[0,1]}$ be a schedule such that $0\leq\eta_{t}\leq\sigma_{t}$.
From the following equality in law which follows from the fact that $\pi_{1}$ is a Gaussian distribution,
$\sigma_{t}X_{1}\overset{\mathcal{L}}{=}(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\tilde{X}_{1}+\big(\sigma_{t}-(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\big)\mu+\eta_{t}\Sigma^{1/2}Z\;,\quad(\tilde{X}_{1},Z)\sim\pi_{1}\otimes\mathcal{N}(0,\mathbf{I}_{d})$
we find that
$X_{t}\overset{\mathcal{L}}{=}\alpha_{t}X_{0}+(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}X_{1}+\big(\sigma_{t}-(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\big)\mu+\eta_{t}\Sigma^{1/2}Z_{t},$ (42)
where $(X_{0},X_{1},Z_{t})\sim\pi_{0}\otimes\pi_{1}\otimes\mathcal{N}(0,\mathbf{I}_{d})$. We thus define for any $t\in[0,1]$ the following bridge transitions,
$q^{\eta}_{\smash{t\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{t}|x_{0},x_{1})\coloneqq\begin{cases}\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t}),&\eta_{t}^{2}=0\;,\\ \mathrm{N}(x_{t};\alpha_{t}x_{0}+(\sigma^{2}_{t}-\eta_{t}^{2})^{1/2}x_{1}+\big(\sigma_{t}-(\sigma^{2}_{t}-\eta^{2}_{t})^{1/2}\big)\mu,\,\eta^{2}_{t}\Sigma)\mathrm{d}x_{t},&\eta_{t}^{2}>0\;.\end{cases}$ (43)
Then clearly from (41) and (42) we have for all $\eta_{t}\in[0,\sigma_{t}]$,
$\pi_{t}(\mathrm{d}x_{t})=\int q^{\eta}_{\smash{t\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{t}|x_{0},x_{1})\,\pi_{0}(x_{0})\pi_{1}(x_{1})\mathrm{d}x_{0}\mathrm{d}x_{1}\;.$ (44)
Now define the reverse transition
$\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{s}|x_{t})\coloneqq\int q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},x_{1})\;\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})$ (45)
where $\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t})$ denotes the conditional law of $(X_{0},X_{1})$ given $X_{t}=x_{t}$ under the joint distribution induced by (41). This conditional can be written as $\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})=\updelta_{(x_{t}-\alpha_{t}x_{0})/\sigma_{t}}(\mathrm{d}x_{1})\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{0}|x_{t}),$
$\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})=\frac{\pi_{0}(x_{0})\,\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)}{\pi_{t}(x_{t})}.$ (46)
Indeed, for any bounded measurable function $f$,
$\int f(x_{0},x_{t},x_{1})\,\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})$ $=\int f(x_{0},x_{t},x_{1})\,\updelta_{\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}}}(\mathrm{d}x_{1})\pi_{0}(x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{t}$
$=\int f(x_{0},x_{t},\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}})\,\pi_{0}(x_{0})\mathrm{N}(x_{t};\alpha_{t}x_{0}+\sigma_{t}\mu,\sigma_{t}^{2}\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{t}$
$=\int f(x_{0},\alpha_{t}x_{0}+\sigma_{t}\mu+\sigma_{t}\Sigma^{1/2}z,\mu+\Sigma^{1/2}z)\,\pi_{0}(x_{0})\mathrm{N}(z;0,\mathbf{I}_{d})\,\mathrm{d}x_{0}\mathrm{d}z$
$=\int f(x_{0},\alpha_{t}x_{0}+\sigma_{t}x_{1},x_{1})\,\pi_{0}(x_{0})\mathrm{N}(x_{1};\mu,\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{1}$
$=\int f(x_{0},x_{t},x_{1})\,\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t})\pi_{0}(x_{0})\mathrm{N}(x_{1};\mu,\Sigma)\mathrm{d}x_{0}\mathrm{d}x_{1}$
which shows that
$\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})=\updelta_{\alpha_{t}x_{0}+\sigma_{t}x_{1}}(\mathrm{d}x_{t})\pi_{0}(x_{0})\pi_{1}(x_{1})\mathrm{d}x_{0}\mathrm{d}x_{1}\;,$ (47)
where the r.h.s. is the joint distribution defined by (41). It then follows that
$\int\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{s}|x_{t})\,\pi_{t}(\mathrm{d}x_{t})$ $=\int q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},x_{1})\,\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})\pi_{t}(\mathrm{d}x_{t})$
$=\int q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},x_{1})\,\pi_{0}(x_{0})\pi_{1}(x_{1})\mathrm{d}x_{0}\mathrm{d}x_{1}$
$=\pi_{s}(x_{s})$
where the second line follows from integrating the r.h.s. in (47) w.r.t. $x_{t}$ and the third one from (44). Finally, by noting that
$\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}x_{s}|x_{t})$ $=\int q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},x_{1})\,\pi_{0,1\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\mathrm{d}(x_{0},x_{1})|x_{t})$
$=\int\underbrace{q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\mathrm{d}x_{s}|x_{0},\frac{x_{t}-\alpha_{t}x_{0}}{\sigma_{t}})}_{q^{\eta}_{\smash{s|0,t}}(\cdot|x_{0},x_{t})}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$
where the defined $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|x_{0},x_{t})$, up to the notation, is exactly the DDIM bridge transition song2021ddim when $\mu=0_{d}$ and $\Sigma=\mathbf{I}_{d}$. Finally, the Gaussian approximation $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|\hat{x}_{0}(x_{t},t),x_{t})$ used at inference, with $\hat{x}_{0}(x_{t},t)\coloneqq\int x_{0}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$, is the one solving
$\operatorname*{argmin}_{r_{s|t}(\cdot|x_{t})\in\mathcal{G}_{\eta^{2}_{s}\Sigma}}\mathsf{KL}(\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t})\|r_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x_{t}))\;,$
where $\mathcal{G}_{\eta^{2}_{s}\Sigma}\coloneqq\{\mathcal{N}(\mu,\eta^{2}_{s}\Sigma):\;\mu\in{\mathbb{R}}^{d}\}$ is the set of Gaussian distributions with covariance set to $\eta^{2}_{s}\Sigma$.
C.2 Explicit denoiser for categorical distributions
In this section we extend the derivation in (3.2) to the case where
$\pi_{1}=\bigotimes_{i=1}^{L}\mathcal{N}(\mu^{i},\mathrm{Diag}(v^{i})),\qquad\mu^{i},v^{i}\in{\mathbb{R}}^{K},\quad v^{ij}>0\;.$
Following (42) and the factorization (2), we still have $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\propto\prod_{i=1}^{L}\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x^{i}_{0}|x^{i}_{t})$
$\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x^{i}_{0}|x^{i}_{t})\;\propto\;\pi^{i}_{\!\theta}(x^{i}_{0})\,\mathrm{N}\big(x^{i}_{t};\,\alpha_{t}x^{i}_{0}+\sigma_{t}\mu^{i},\;\sigma_{t}^{2}\mathrm{Diag}(v^{i})\big)\;.$ (48)
With this structure, the denoiser $\hat{x}^{\theta}_{0}(x_{t},t)\coloneqq\sum_{x_{0}}x_{0}\,\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})$ simplifies to a matrix of posterior probabilities due to the one-hot structure; i.e. for any $i\in[L]$ and $j\in[K]$, $\hat{x}^{\theta}_{0}(x_{t},t)^{ij}=\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(e_{j}|x_{t}).$ Using that
$\mathrm{N}(x^{i}_{t};\alpha_{t}e_{j}+\sigma_{t}\mu^{i},\sigma_{t}^{2}\mathrm{Diag}(v^{i}))\propto\exp\!\left(-\frac{1}{2\sigma_{t}^{2}}\sum_{k=1}^{K}\frac{(x^{ik}_{t}-\alpha_{t}e_{j}^{k}-\sigma_{t}\mu^{ik})^{2}}{v^{ik}}\right),$
we expand the quadratic term and drop all terms independent of $j$ to obtain the logits
$\log\hat{x}^{\theta}_{0}(x_{t},t)^{ij}=\log\varphi^{ij}_{\theta}+\frac{\alpha_{t}}{\sigma_{t}^{2}}\frac{x_{t}^{ij}-\sigma_{t}\mu^{ij}}{v^{ij}}-\frac{\alpha_{t}^{2}}{2\sigma_{t}^{2}}\frac{1}{v^{ij}}+C(i,t)\;.$ (49)
Equivalently, for each $(i,j)\in[L]\times[K]$,
$\hat{x}^{\theta}_{0}(x_{t},t)^{ij}=\frac{\pi^{i}_{\!\theta}(e_{j})\exp\big(\frac{\alpha_{t}}{\sigma^{2}_{t}v^{ij}}(x_{t}-\sigma_{t}\mu^{ij}-\frac{\alpha_{t}}{2})\big)}{\sum_{k=1}^{K}\pi^{i}_{\!\theta}(e_{k})\exp\big(\frac{\alpha_{t}}{\sigma^{2}_{t}v^{ik}}(x_{t}-\sigma_{t}\mu^{ik}-\frac{\alpha_{t}}{2})\big)}$ (50)
which yields $\hat{x}^{\theta}_{0}(x_{t},t)\!=\!\text{softmax}(\varphi_{\theta}+\frac{\alpha_{t}\lambda}{\sigma^{2}_{t}}\odot(x_{t}-\sigma_{t}\mu-\frac{\alpha_{t}}{2}\mathbf{1})).$ where $\lambda\in{\mathbb{R}}^{L\times K}$ with $\lambda^{i,j}=1/v^{i,j}$ and $\mathbf{1}\in{\mathbb{R}}^{L\times K}$ is the all-ones matrix.
Appendix D Experimental Details
D.1 Baselines
We compare our method against three representative baselines from the literature: the Straight-Through (ST) estimator (bengio2012estimating), the Gumbel-Softmax estimator (more precisely, its Straight-Through variant) (jang2017categorical), and the more recent ReinMax method (liu2023reinmax). Among these, ReinMax reports state-of-the-art performance on most of the benchmarks it considers and, to the best of our knowledge, is one of the most recent approaches addressing the same class of problems as ours. For this reason, and since liu2023reinmax show that ReinMax consistently outperforms several earlier alternatives, we do not include additional baselines in our comparison. For the hyperparameters of the other samplers, Gumbel-Softmax and ReinMax, the only one to tune is the temperature $\tau$ and we choose values similar to those used in (jang2017categorical; liu2023reinmax), noting that as underlined in (liu2023reinmax), ReinMax works better with moderate or higher $\tau$ whereas, Gumbel-Softmax works better with lower $\tau$.
D.2 Implementation and hyperparameters
For both the polynomial programming experiment and the categorical VAE, we closely follow the experimental settings of liu2023reinmax, which themselves build upon fan2022trainingdiscretedeepgenerative and earlier studies. To ensure consistency across experiments and to limit the risk of overfitting, we use the same optimizer (Adam) for all methods and keep the hyperparameters as similar as possible whenever this is meaningful.
Table 5 summarizes the main implementation details and hyperparameters for each sampler. For each method, we report the optimizer, learning rate, and sampler-specific parameters.
Table 5: Hyperparameters used for each experiment. For each benchmark, we report the optimizer, learning rate, and sampler-specific parameters for every sampler.
ReDGE ReDGE-Max ReDGE-Cov Gumbel-Softmax Straight-Through ReinMax
Polynomial Programming Optimizer Adam Adam Adam Adam Adam Adam
Learning rate 0.05 0.05 0.05 0.05 0.05 0.05
Sampler params n=7n{=}7 $n{=}10,{t_{1}}{=}0.7$ n=5n{=}5 $\tau{=}0.1$ $\tau{=}1.0$ $\tau{=}1.0$
GMM Variational Inference Optimizer Adam Adam Adam Adam Adam Adam
Learning rate 0.01 0.01 0.01 0.01 0.01 0.01
Sampler params n=4n{=}4 $n{=}10,{t_{1}}{=}0.5$ n=3n{=}3 $\tau{=}0.1$ $\tau{=}1.0$ $\tau{=}1.0$
Sudoku Optimizer Adam Adam Adam Adam Adam Adam
Learning rate 0.05 0.05 0.05 0.05 0.05 0.05
Sampler params n=4n{=}4 $n{=}10,{t_{1}}{=}0.5$ n=3n{=}3 $\tau{=}0.3$ $\tau{=}1.0$ $\tau{=}1.0$
Categorical VAE Optimizer Adam Adam Adam Adam Adam Adam
Learning rate 5e-4 5e-4 5e-4 5e-4 5e-4 5e-4
Sampler params n=4n{=}4 $n{=}10,{t_{1}}{=}0.5$ n=3n{=}3 $\tau{=}0.5$ $\tau{=}1.0$ $\tau{=}1.0$
Polynomial Programming
Following (liu2023reinmax), we use a batch size of 256, a length of 128, 2 categorical dimensions and a vector $c:=(c_{1},\dots,c_{L})\in\mathbb{R}^{L}$, $\forall i,c_{i}=0.45$.
GMM
For the GMM experiment, we use a random initialization of the clutering parameters as well as of the means. We take the follwowing hyperparameters: $D=2$, $K=20$, $\sigma_{0}=15$, $\sigma_{y}=2$ and $500$ samples.
Categorical VAE
Following prior work (liu2023reinmax; fan2022trainingdiscretedeepgenerative), we use two-layer multilayer perceptrons (MLPs) for both the encoder and the decoder. The encoder consists of hidden layers of sizes 512 and 256, while the decoder uses hidden layers of sizes 256 and 512. We set the batch size to 200, use LeakyReLU activations, and train for $200$ epochs, which corresponds to $60000$ optimization steps.
Sudoku
We use the Ritvik19/Sudoku-Dataset from HuggingFace and run 1000 optimization steps per sudoku. We test the method on the first 2000 sudokus of the training set.
D.3 Additional Results
Polynomial Programming
As mentioned in 4.1, using a linear relaxation for the polynomial programming problem yield a very different and simpler optimization problem, which is nevertheless exactly the same in terms of its optimum.
We see in Figure 4 that in this setting Straight-Through achieves the best performance.
(a) $p=1.5$
(b) $p=2.0$
(c) $p=3.0$
Figure 4: Polynomial programming benchmark for different values of the exponent $p$ with the linear relaxation.
GMM
We show the negative ELBO curves for 3 differents seeds over the optimization in Fig. 5.
Figure 5: Negative ELBO curves during optimization for the GMM experiment.
Categorical VAE
We show the loss curves of the two other Categorical VAE experiment configurations in Fig 6. We see that in all the cases, not only does ReDGE-Cov achieve better final performance, it also converges faster and in a more stable way.
(a) L=32,K=64L=32,\,K=64
(b) L=30,K=10L=30,\,K=10
Figure 6: Categorical VAE training curves for two additional latent–categorical configurations.